### 31 Mar 2026 v1

# Stochastic Cahn–Hilliard Equations from One-Dimensional Ising–Kac–Kawasaki Dynamics

Abstract
We study the scaling limit of the one-dimensional lattice Ising–Kac–Kawasaki dynamics under conservative Kawasaki exchange rate. For the Kac coarse-grained field $X_{\gamma}$, we derive a martingale formulation with a discrete conservative drift and a Dynkin martingale. The nonlinear drift is identified by a conservative multiscale replacement scheme based on one-block/two-block estimates, yielding a cubic conservative term in the macroscopic equation. For the stochastic part, we compute the predictable quadratic variation, and obtain a divergence-form Gaussian noise. As a consequence, $X_{\gamma}$ converges to a one-dimensional stochastic Cahn–Hilliard equation with conserved noise.
Keywords: Ising-Kac model; Kawasaki dynamics; exclusion process; stochastic Cahn–Hilliard equation; hydrodynamic limit.
1 Introduction
Phase separation with a conserved order parameter is classically described at the continuum level by the Cahn–Hilliard equation, or equivalently by Model B in the Hohenberg–Halperin classification [1, 13, 19]. Its stochastic counterpart is expected to govern mesoscopic fluctuations around phase-separating profiles under a conservation law, and therefore naturally takes the form of a conservative SPDE, namely a stochastic Cahn–Hilliard equation driven by divergence-type noise. In its deterministic form, it can be interpreted as an $H^{-1}$-gradient flow of a Ginzburg–Landau-type free energy, which naturally enforces mass conservation. Random perturbations—motivated by thermal fluctuations and mesoscopic coarse-graining—lead to the stochastic Cahn–Hilliard equation, typically driven by a divergence-form noise so that the conservation law is preserved at the stochastic level [8]. A central problem in nonequilibrium statistical mechanics is to derive such a macroscopic conservative SPDE directly from a microscopic conservative interacting particle system [21].
In this work we consider the one-dimensional Ising–Kac model [14] under Kawasaki exchange dynamics on $\Lambda_{N}:=\mathbb{Z}/(2N+1)\mathbb{Z}$. Let $\Sigma_{N}:=\{-1,1\}^{\Lambda_{N}}.$ We consider a kernel $\kappa_{\gamma}(z)$ of the form with a Kac parameter $\gamma\in(0,\frac{1}{4})$,
$\kappa_{\gamma}(z)=\gamma\,\mathfrak{K}(\gamma z),\quad\kappa_{\gamma}(z)\geq 0,\quad\sum_{z\in\Lambda_{N}}\kappa_{\gamma}(z)=1,$
and the Ising–Kac Hamiltonian
$H_{\gamma}(\sigma)=-\frac{1}{2}\sum_{i,j\in\Lambda_{N}}\kappa_{\gamma}(i-j)\,\sigma_{i}\sigma_{j},\quad\sigma\in\Sigma_{N}.$
The Ising–Kac–Kawasaki dynamics $\sigma(t)$ is a continuous-time Markov chain on $\Sigma_{N}$ generated by nearest-neighbour exchanges: for each bond $(i,i+1)$ let $\sigma^{i,i+1}$ denote the configuration obtained by swapping $\sigma_{i}$ and $\sigma_{i+1}$, and define the heat-bath exchange rate
$c_{i}(\sigma):=\frac{1}{1+\exp\!\bigl(\beta\,\Delta_{i}H_{\gamma}(\sigma)\bigr)},\qquad\Delta_{i}H_{\gamma}(\sigma):=H_{\gamma}(\sigma^{i,i+1})-H_{\gamma}(\sigma).$
The generator is
$(\mathscr{L}_{\gamma}^{N}f)(\sigma)=\sum_{i\in\Lambda_{N}}c_{i}(\sigma)\,\bigl[f(\sigma^{i,i+1})-f(\sigma)\bigr].$
A key structural feature is magnetization conservation: we define the oriented microscopic current associated with bond $(i,i+1)$ as
$j_{i}(\sigma):=(\sigma_{i}-\sigma_{i+1})\,c_{i}(\sigma),$
then a direct computation yields the discrete continuity equation
$(\mathscr{L}_{\gamma}^{N}\sigma_{i})(\sigma)=j_{i-1}(\sigma)-j_{i}(\sigma),$
i.e. the change at site $i$ is “incoming current minus outgoing current”. Hence the spin variation at site $i$ is given by the incoming current through the left bond minus the outgoing current through the right bond, which is the natural lattice analogue of a divergence form.
For a fixed total magnetization $M(N):=\sum_{i\in\Lambda_{N}}\sigma_{i}$, we define the canonical sector
$\Sigma_{N,M}:=\left\{\sigma\in\Sigma_{N}:\ \sum_{i\in\Lambda_{N}}\sigma_{i}=M(N)\right\}.$
The corresponding canonical Gibbs measure is
$\mu_{N,\gamma,\beta}(d\sigma):=\frac{1}{Z_{N,\gamma,\beta}^{M}}\exp\!\bigl(-\beta H_{\gamma}(\sigma)\bigr)\mathbf{1}_{\Sigma_{N,M}}(\sigma),$
where $Z_{N,\gamma,\beta}^{M}:=\sum_{\sigma\in\Sigma_{N,M}}$ is the normalizing constant.
Since the Kawasaki exchange $\sigma\mapsto\sigma^{i,i+1}$ preserves the total magnetization, the dynamics leaves each sector $\Sigma_{N,M}$ invariant. Moreover, the exchange rates $c_{i}(\sigma)$ satisfy detailed balance with respect to $\mu_{N,\gamma,\beta}$. Hence the Ising–Kac–Kawasaki dynamics is reversible on each canonical sector:
$\sum_{i\in\Lambda_{N}}\sigma_{i}(t)=\sum_{i\in\Lambda_{N}}\sigma_{i}(0)=M(N)\qquad\text{for all }t\geq 0.$
To connect to a macroscopic field, we introduce the Kac-smoothed spin field
$h_{\gamma}(\sigma;i):=\sum_{j\in\Lambda_{N}}\kappa_{\gamma}(i-j)\sigma_{j},\quad i\in\Lambda_{N},$
and define the the Kac coarse-grained field $X_{\gamma}(t,\varepsilon_{\gamma}i)=\delta_{\gamma}^{-1}h_{\gamma}\!\left(\sigma\big(\frac{t}{\alpha_{\gamma}}\big);i\right),$ where $\varepsilon_{\gamma}i\in\mathbb{T}$, and $(\alpha_{\gamma},\varepsilon_{\gamma},\delta_{\gamma})$ are the time, space and amplitude scales. Since the Kac kernel is normalized by $\sum_{z\in\Lambda_{N}}\kappa_{\gamma}(z)=1,$ we have $\sum_{i\in\Lambda_{N}}h_{\gamma}(\sigma;i)=\sum_{j\in\Lambda_{N}}\sigma_{j}=M$. Thus for the Kac coarse-grained field $X_{\gamma}$ also has conservation law:
$\langle X_{\gamma}(t),1\rangle_{\gamma}=\langle X_{\gamma}(0),1\rangle_{\gamma}=\frac{\varepsilon_{\gamma}}{\delta_{\gamma}}M(N)\qquad\text{for all }t\in[0,T].$
For a smooth test function $\phi\in C^{\infty}(\mathbb{T})$, set
$\langle X_{\gamma}(t),\phi\rangle_{\gamma}:=\varepsilon_{\gamma}\sum_{i\in\Lambda_{N}}X_{\gamma}(t,\varepsilon_{\gamma}i)\phi(\varepsilon_{\gamma}i).$
Using the Dynkin martingale decomposition, we have the canonical drift–martingale splitting:
$\langle X_{\gamma}(t),\phi\rangle_{\gamma}=\langle X_{\gamma}(0),\phi\rangle_{\gamma}+\int_{0}^{t}\Bigl(\alpha_{\gamma}^{-1}\mathscr{L}_{\gamma}^{N}\langle X_{\gamma},\phi\rangle_{\gamma}\Bigr)(\sigma_{s/\alpha_{\gamma}})\,ds+M_{t}^{\gamma}(\phi),$ (1)
where $M_{t}^{\gamma}(\phi)$ is a Dynkin martingale.
This formula already contains the full structure of the limit problem: the first integral term is the discrete conservative drift, while $M^{\gamma}$ is the microscopic noise source. Under the scaling considered in this work, one expects the limiting dynamics is a one-dimensional stochastic Cahn–Hilliard equation of the form
$\partial_{t}X=-\nu\Delta^{2}X-A\Delta X+\chi\Delta(X^{3})+\sigma_{*}\nabla\cdot\xi,\quad\int_{\mathbb{T}}X(\cdot,x)dx\equiv M.$ (2)
Here $\xi$ is space–time white noise and $\partial_{x}\xi$ is the conservative noise forcing. The coefficients $(\nu,A,\chi,\sigma_{*})$ are determined by the Kac kernel, the inverse temperature, and the scaling relations between $(\alpha_{\gamma},\varepsilon_{\gamma},\delta_{\gamma})$. Here, the fourth-order term is generated by the Kac interaction after expansion of the coarse-grained field, the second-order linear term comes from the critical tuning, the cubic conservative drift comes from the nonlinear closure of the microscopic current, and the conservative noise $\partial_{x}\xi$ arises from the martingale part of the microscopic dynamics.
From the PDE/SPDE viewpoint, stochastic Cahn–Hilliard equations provide a rich testing ground where high-order dissipation, conservation constraints, and nonlinear potentials interact. Early foundational work developed a semigroup/variational framework and energy estimates for stochastic Cahn–Hilliard-type dynamics [4, 3]. More recent advances address physically relevant features such as conservative (divergence-type) noise, singular/logarithmic potentials enforcing physical constraints, and less restrictive growth assumptions on noise coefficients; see, for instance, [2, 5, 9]. On the long-time side, ergodicity under mass conservation and singular potentials has been investigated via probabilistic tools such as (log-)Harnack inequalities [9], while random dynamical systems methods provide an abstract and robust route to random attractors for broad classes of (locally monotone) SPDEs encompassing Cahn–Hilliard-type models [7]. In particular, conservative stochastic Cahn–Hilliard equations in two dimensions and related invariant measures are considered in [20]. These results provide a robust macroscopic analytic framework, but they do not by themselves explain how the conservative SPDE emerges from a microscopic lattice dynamics.
On the interacting-particle side, hydrodynamic and fluctuation theories for conservative systems have developed a powerful multiscale toolbox based on local equilibrium, the Boltzmann–Gibbs principle, spectral-gap estimates, and one-block/two-block replacements [21, 15, 6]. These methods explain how microscopic conservative exchanges generate macroscopic conservation laws, and they strongly suggest that the right route to stochastic Model B is through a careful closure of the microscopic current and of the local mobility. Beyond weakly interacting settings, quantitative hydrodynamic limits for conservative dynamics with strong, finite-range interactions have also become accessible [17], highlighting the reach of modern two-scale and functional inequality approaches. However, applying this philosophy to the critical Ising–Kac–Kawasaki regime is nontrivial because both the nonlinear drift and the noise must be extracted from conservative bond observables.
It is also useful to compare the present conservative setting with recent discrete-to-SPDE results for non-conservative (Ising–Kac–Glauber / Model A) dynamics. Mourrat and Weber proved that the two-dimensional dynamical Ising–Kac model converges to the renormalized $\Phi^{4}_{2}$ equation [18], and Grazieschi, Matetski and Weber obtained the corresponding three-dimensional convergence to $\Phi^{4}_{3}$ [10]. Zhu and Zhu established lattice approximation results for the dynamical $\Phi^{4}_{3}$ model [23], while Gubinelli and Hofmanová constructed the Euclidean $\Phi^{4}_{3}$ field by PDE/Dirichlet-form methods [11]. These works show that renormalized singular SPDEs can indeed arise as scaling limits of discrete systems, and that lattice-level counterterms reflect the structure of the continuum limit. By contrast, the conservative case requires an additional current-level analysis: one must close a divergence-form nonlinear drift and a divergence-form noise simultaneously.
Motivated by these developments, a central challenge remains to rigorously connect lattice Ising–Kac–Kawasaki dynamics near criticality with macroscopic SPDEs, in a way that simultaneously identifies the nonlinear conservative drift and the divergence-form noise from the microscopic generator and its martingale part. In this paper we aim to advance the lattice-to-SPDE direction and, building on the same multi-scale replacement philosophy, to push beyond the dynamical limit toward the associated equilibrium and long-time theory: convergence of discrete (canonical) Gibbs measures to the macroscopic constrained $\Phi^{4}$-type measures, and the emergence of reversibility and ergodic behavior in the scaling limit.
However, deriving this stochastic Cahn–Hilliard equation rigorously from an underlying conservative lattice dynamics remains a delicate problem: the nonlinearity in the microscopic bond current must be closed in terms of the macroscopic field, and the martingale fluctuations must be identified as a divergence-form Gaussian noise. Both the nonlinear drift term and the noise term must be derived from microscopic currents.
The drift term in (1) can be rewritten in conservative form using the microscopic bond current. A direct computation yields an explicit expression for the energy difference $\Delta_{x}H_{\gamma}$ across a swap, which in turn permits a Taylor expansion of the heat-bath rate. This produces a leading nonlinear current term containing the fast local observable $d^{2}_{x}:=(\sigma_{x}-\sigma_{x+1})^{2}$, which is not itself a function of the coarse-grained field and must be closed by a conservative multiscale replacement argument. This yields a closed macroscopic current of the form $X^{2}\nabla X$, and after taking the outer discrete divergence one obtains the cubic conservative drift $\partial_{x}^{2}(X^{3})$.
Moreover, the quadratic variation of the discrete martingales $M^{\gamma}(\phi)$ involves the local mobility observable $q_{i}:=d^{2}_{i}c_{i}(\sigma)=(\sigma_{i}-\sigma_{i+1})^{2}c_{i}(\sigma),$ and this must be closed by the same conservative replacement mechanism in order to produce a deterministic limiting covariance of divergence type. A martingale central limit theorem then yields convergence to a Gaussian limit with covariance $\int_{\mathbb{T}}|\partial_{x}\phi|^{2}dx$, which identifies the noise as the conservative forcing $\partial_{x}\xi$.
In this paper we address these issues for the one-dimensional Ising–Kac model evolving under Kawasaki exchange dynamics. Our goal is to show that, under a suitable scaling, a Kac coarse-grained field converges to a one-dimensional stochastic Cahn–Hilliard equation.
Assumptions and main result.
Throughout out this paper, assume that following assumptions hold.
  1. (A)
    (Kernel assumptions) The Kac kernel $\kappa_{\gamma}(z)$ is of the form $\kappa_{\gamma}(z)=\gamma\,\mathfrak{K}(\gamma z),\quad\kappa_{\gamma}(z)=\kappa_{\gamma}(-z),$ where the positive function $\mathfrak{K}\in W^{1,1}(\mathbb{R})\cap W^{1,\infty}(\mathbb{R}),$ and
    $\sum_{z\in\Lambda_{N}}\kappa_{\gamma}(z)=1,\quad\mathfrak{m}_{2}:=\int_{\mathbb{R}}u^{2}\,\mathfrak{K}(u)\,du<\infty.$
  2. (B)
    (Scaling assumptions) There exist positive scale factors $\alpha_{\gamma},\delta_{\gamma},\varepsilon_{\gamma}$ such that
    $\varepsilon_{\gamma}:=\frac{1}{N}\to 0,\qquad\frac{\varepsilon_{\gamma}}{\gamma}\to 0,\qquad\frac{\varepsilon_{\gamma}^{2}}{\delta_{\gamma}}\to 0,\qquad\frac{\varepsilon_{\gamma}^{3}}{\alpha_{\gamma}}\to 0,\qquad\gamma\,\frac{\varepsilon_{\gamma}^{3}}{\alpha_{\gamma}\delta_{\gamma}^{2}}\to 0.$
    We also assume that
    $\frac{\mathfrak{m}_{2}}{4}\frac{\varepsilon_{\gamma}^{4}}{\alpha_{\gamma}\gamma^{2}}\to\nu>0,\quad\lambda_{\gamma}:=\frac{\varepsilon_{\gamma}^{2}}{\alpha_{\gamma}}\to\lambda>0,\quad\frac{\varepsilon_{\gamma}^{3}}{\alpha_{\gamma}\delta_{\gamma}^{2}}\rightarrow\sigma_{*}^{2}>0.$
  3. (C)
    (Initial law and entropy bound) Let $f_{N,0}\mu_{N,\gamma,\beta}$ be the initial law, where $\mu_{N,\gamma,\beta}$ is the Gibbs measure associated with $H_{\gamma}$. We assume
    $H_{N}(f_{N,0}\mid\mu_{N,\gamma,\beta})\leq C_{H}N,$
    and as $\gamma\to 0$, the initial data converges in the sense of
    $\langle X_{\gamma}(0),1\rangle_{\gamma}\to\langle X_{\gamma}(0),1\rangle:=M,\quad\widetilde{X}_{\gamma}(0)\longrightarrow X_{0}\quad\text{in }H^{-1}(\mathbb{T}).$
The following main result of the present paper is that the Kac-smoothed field $X_{\gamma}$ converges, in the sense of the associated martingale problem, to the one-dimensional stochastic Cahn–Hilliard / Model B equation with conservative noise. This theorem is a unified conservative replacement scheme that closes the nonlinear current and the mobility in parallel, leading to a complete martingale-problem identification of the 1D stochastic Cahn–Hilliard limit equations.
Theorem 1.
Assume that the assumption $({\bf A})$-$({\bf C})$ hold. After Fourier extension and letting $\gamma\to 0$, the Kac coarse-grained field $X_{\gamma}$ converges in law to $X$ in $L^{2}(0,T;L^{2}(\mathbb{T}))\cap C([0,T],H^{-3}(\mathbb{T}))$, where $X$ is the unique solution of the one-dimensional stochastic Cahn–Hilliard equation
$\partial_{t}X=-\nu\Delta^{2}X-\frac{(1+\beta)\lambda}{2}\Delta X+\frac{\lambda\beta}{6}\Delta(X^{3})+\sigma_{*}\nabla\cdot\xi.$ (3)
Moreover, $X(\cdot)$ has conservation law:
$\langle X(t),1\rangle=\langle X(0),1\rangle=M\qquad\text{for all }t\in[0,T].$
The rest of this paper is organized as follows: Section 2 introduces the Ising–Kac–Kawasaki dynamics, the Dirichlet forms, and basic well-posedness results of one-dimensional stochastic Cahn–Hilliard equation. Section 3 rewrites the drift in current form and performs the Taylor expansion leading to the nonlinear current term. Section 4 establish one-block and two-block estimates and derive the second-order current closure. Section 5 computes the quadratic variation of the Dynkin martingale, proves the mobility replacement, and identifies the conservative Gaussian noise. We establish the discrete $H^{-1}$ energy estimates for $X_{\gamma}$ in Section 6 . Finally, Section 7 combines tightness and identification to conclude the proof of the main theorem.
2 Preliminary
2.1 Fourier extension from the lattice to the torus
To place the whole argument in a fixed function space over $\mathbb{T}$, we need identify lattice fields with trigonometric polynomials via discrete Fourier interpolation, following the Fourier-model viewpoint used in weakly nonlinear lattice-to-SPDE arguments.
Let $\Lambda_{\varepsilon_{\gamma}}:=\varepsilon_{\gamma}\mathbb{Z}/\mathbb{Z}=\Bigl\{-\frac{1}{2},-\frac{1}{2}+\varepsilon_{\gamma},\dots,-\varepsilon_{\gamma},0,\varepsilon_{\gamma},\dots,\frac{1}{2}\Bigr\}.$ We also define the finite-dimensional trigonometric polynomial space
$\mathcal{T}_{\gamma}:=\mathrm{span}\{e_{k}(x):=e^{2\pi ikx},\ k\in\mathbb{Z}_{\gamma}\}\subset C^{\infty}(\mathbb{T}).$
For a lattice function $f:\Lambda_{\varepsilon_{\gamma}}\to\mathbb{C}$, define its discrete Fourier coefficients by
$\widehat{f}_{\gamma}(k):=\varepsilon_{\gamma}\sum_{j\in\Lambda_{\varepsilon_{\gamma}}}f(j)e^{-2\pi ikj},\quad k\in\mathcal{B}_{\gamma}.$
where $\mathcal{B}_{\gamma}:=\{k\in\mathbb{Z}:\ |k|\leq c\,\gamma^{-1}\}$ is the Brillouin zone. Its Fourier extension to $\mathbb{T}$ is
$(\mathrm{Ext}_{\gamma}f)(x):=\sum_{k\in\mathcal{B}_{\gamma}}\widehat{f}_{\gamma}(k)e_{k}(x),\quad e_{k}(x):=e^{2\pi ikx}.$
Let $x_{i}:=\varepsilon_{\gamma}i\in\mathbb{T}$, and define the discrete gradient and Laplacian as
$\nabla_{\varepsilon_{\gamma}}f(x_{i}):=\frac{f(x_{i+1})-f(x_{i})}{\varepsilon_{\gamma}},\quad\Delta_{\varepsilon_{\gamma}}:=\nabla_{\varepsilon_{\gamma}}^{*}\nabla_{\varepsilon_{\gamma}},\quad f\in\mathcal{T}_{\gamma}.$
We also set $A_{\gamma}:=-\Delta_{\varepsilon_{\gamma}},$ Define the extended discrete Laplacian $\Delta_{\varepsilon_{\gamma}}$ on $\mathcal{T}_{\gamma}$ by Fourier multiplier:
$\Delta_{\varepsilon_{\gamma}}e_{k}=-\lambda_{\gamma}(k)e_{k},\quad\lambda_{\gamma}(k):=\frac{4}{\varepsilon_{\gamma}^{2}}\sin^{2}(\pi k\varepsilon_{\gamma}),\qquad k\in\mathcal{B}_{\gamma}.$
This is the torus realization of $A_{\gamma}=-\Delta_{\varepsilon_{\gamma}}$.
Now we show the symbol comparison of $A_{\gamma}=-\Delta_{\varepsilon_{\gamma}}$ and second-order consistency.
Lemma 1.
Let $\lambda(k):=(2\pi k)^{4}.$ Then for every $k\in\mathcal{B}_{\gamma}$ one has
$c\,|k|^{4}\leq\lambda_{\gamma}(k)\leq C\,|k|^{4},\quad\text{and}\quad|\lambda_{\gamma}(k)-\lambda(k)|\leq C\,\varepsilon_{\gamma}^{2}|k|^{6}.$
where the constants $c,C>0$ are independent of $\varepsilon_{\gamma}$ and $k$.
Proof.
For $|k|\leq N=(\varepsilon_{\gamma})^{-1}$ we have $|\pi\varepsilon k|\leq\pi/2$. Hence
$\frac{2}{\pi}|y|\leq|\sin y|\leq|y|,\qquad|y|\leq\frac{\pi}{2}.$
With $y=\pi\varepsilon_{\gamma}k$, this yields
$16\varepsilon_{\gamma}^{-4}\Bigl(\frac{2}{\pi}\pi\varepsilon_{\gamma}|k|\Bigr)^{4}\leq\lambda_{\varepsilon_{\gamma}}(k)\leq 16\varepsilon^{-4}(\pi\varepsilon_{\gamma}|k|)^{4},$
which is exactly
$c|k|^{4}\leq\lambda_{\varepsilon}(k)\leq C|k|^{4}.$
For the consistency estimate, note that
$\sin y=y-\frac{y^{3}}{6}+O(y^{5}),\qquad y\to 0.$
Therefore
$\lambda_{\gamma}(k)=16\varepsilon_{\gamma}^{-4}\bigl((\pi\varepsilon_{\gamma}k)^{4}+O((\pi\varepsilon_{\gamma}k)^{6})\bigr)=(2\pi k)^{4}+O(\varepsilon_{\gamma}^{2}|k|^{6}).$
This proves the claim. ∎
Now we show the compatibility of the extension with the discrete Laplacian.
Lemma 2.
For every lattice field $u$ and every polynomial $P$,
$\mathrm{Ext}_{\gamma}(P(A_{\gamma})u)=P(\mathcal{A}_{\gamma})\mathrm{Ext}_{\gamma}u.$
Proof.
The operator $A_{\gamma}$ is diagonal in the discrete Fourier basis, with symbol $\lambda_{\gamma}(k)$. Since $\mathrm{Ext}_{\gamma}$ preserves discrete Fourier coefficients, the equality $\mathrm{Ext}_{\gamma}(A_{\gamma}u)=\mathcal{A}_{\gamma}\mathrm{Ext}_{\gamma}u$ holds mode. The polynomial functional calculus then follows immediately. ∎
For $U=\sum_{k\in\mathbb{Z}_{\gamma}}\widehat{U}(k)e_{k}\in\mathcal{T}_{\gamma}$, define the Sobolev norms
$\|U\|_{H^{s}(\mathbb{T})}^{2}:=\sum_{k\in\mathbb{Z}_{\gamma}}(1+|k|^{2})^{s}|\widehat{U}(k)|^{2}.$
On the lattice side, define the discrete Sobolev norms by
$\|u\|_{H^{s}_{\gamma}}^{2}:=\sum_{k\in\mathbb{Z}_{\gamma}}(1+\lambda_{\gamma}(k))^{s}|\widehat{u}_{\gamma}(k)|^{2}.$
Let $(\varrho_{q})_{q\geq-1}$ be a standard smooth dyadic partition of unity on $\mathbb{Z}$, and let
$\Delta_{j}u:=\sum_{k\in\mathbb{Z}}\varrho_{q}(k)\widehat{u}(k)e_{k}$
be the Littlewood–Paley blocks on $\mathbb{T}$. We define the Besov norm on $\mathbb{T}$ by
$\|u\|_{B^{\beta}_{p,q}}^{p}:=\left(\sum_{j\geq-1}2^{\beta qj}\|\Delta_{j}u\|_{L^{p}(\mathbb{T})}^{q}\right)^{1/q}..$
For $\beta\in\mathbb{R}$ and $1\leq p,q<\infty$, define the discrete Besov norm by
$\|f\|_{B^{\beta}_{p,q,\gamma}}:=\|\mathrm{Ext}_{\gamma}f\|_{B^{\beta}_{p,q}(\mathbb{T})}=\left(\sum_{j\geq-1}2^{\beta qj}\|\Delta_{j}\mathrm{Ext}_{\gamma}f\|_{L^{p}(\mathbb{T})}^{q}\right)^{1/q}.$
Lemma 3 (Equivalence of discrete and continuum Sobolev norms).
For every fixed $s\in[-2,2]$, there exist constants $c_{s},C_{s}>0$, independent of $\gamma$, such that for every lattice field $u$,
$c_{s}\,\|u\|_{H^{s}_{\gamma}}\leq\|\mathrm{Ext}_{\gamma}u\|_{H^{s}(\mathbb{T})}\leq C_{s}\,\|u\|_{H^{s}_{\gamma}}.$
In particular,
$\|\mathrm{Ext}_{\gamma}u\|_{H^{-1}(\mathbb{T})}\asymp\|u\|_{-1,\gamma},\qquad\|\mathrm{Ext}_{\gamma}u\|_{H^{1}(\mathbb{T})}\asymp\|u\|_{1,\gamma}.$
Now we show the approximation and block-wise estimates of semigroup.
Lemma 4.
Let $\lim_{\gamma\to 0}\nu_{\gamma}=\nu$, $S_{\gamma}(t):=e^{-t\nu_{\gamma}\Delta_{\gamma}^{2}}$, and $S(t):=e^{-t\nu\Delta^{2}}.$ Under Assumption $({\bf A})$, the following hold.
  1. (i)
    For every fixed $K>0$,
    $\sup_{|k|\leq K}\sup_{0\leq t\leq T}\bigl|e^{-t\lambda_{\gamma}(k)}-e^{-t(2\pi k)^{4}}\bigr|\longrightarrow 0.$
  2. (ii)
    For every dyadic block $q\geq-1$,
    $\sup_{\gamma}\sup_{k\in\mathrm{supp}\,\varrho_{q}\cap\mathcal{B}_{\gamma}}\int_{0}^{T}e^{-2(T-r)\lambda_{\gamma}(k)}|k|^{2}\,dr\lesssim 2^{-q}.$
  3. (iii)
    There exists $\vartheta\in(0,\frac{1}{4})$ such that for every $q\geq-1$,
    $\sup_{\gamma}\sup_{k\in\mathrm{supp}\,\varrho_{q}\cap\mathcal{B}_{\gamma}}\int_{0}^{s}\bigl|e^{-(t-r)\lambda_{\gamma}(k)}-e^{-(s-r)\lambda_{\gamma}(k)}\bigr|^{2}|k|^{2}\,dr\leq C|t-s|^{\vartheta}2^{-q(2-4\vartheta)},$
    and
    $\sup_{\gamma}\sup_{k\in\mathrm{supp}\,\varrho_{q}\cap\mathcal{B}_{\gamma}}\int_{s}^{t}e^{-2(t-r)\lambda_{\gamma}(k)}|k|^{2}\,dr\leq C|t-s|^{\vartheta}2^{-q(2-4\vartheta)}.$
Proof.
Item (i) follows immediately from the convergence of the symbols $\lambda_{\gamma}(k)\to(2\pi k)^{4}$ on compact frequency sets. Item (ii) follows from
$\int_{0}^{T}e^{-2(T-r)\lambda_{\gamma}(k)}|k|^{2}\,dr\lesssim\frac{|k|^{2}}{\lambda_{\gamma}(k)}\lesssim|k|^{-2},$
and the fact that $|k|\sim 2^{q}$ on the support of $\varrho_{q}$. Item (iii) is the standard semigroup increment estimate: for the first integral, one writes
$e^{-(t-r)\lambda}-e^{-(s-r)\lambda}=e^{-(s-r)\lambda}(e^{-(t-s)\lambda}-1)$
and uses $|e^{-a}-1|\lesssim a^{\vartheta}$ for $a\geq 0$; the second follows by direct integration. ∎
2.2 Well-posedness of the 11d stochastic Cahn–Hilliard equation
We briefly recall the well-posedness result for the 11d stochastic Cahn–Hilliard equation
$\partial_{t}X=-\nu\Delta^{2}X-A\Delta X+\chi\Delta(X^{3})+\sigma_{*}\nabla\cdot\xi.$ (4)
where $\xi$ is a space-time white noise on $\mathbb{R}^{+}\times\mathbb{T}$, $\nu>0$, $\chi>0$, $\sigma_{*}>0$, and $A\in\mathbb{R}$. In one space dimension, this equation is much better behaved than in higher dimensions: the linear stochastic convolution is function-valued, so that the cubic term can be treated as a classical nonlinear term, and one does not need renormalization.
In one space dimension, the stochastic Cahn–Hilliard equation (4)driven by additive conservative noise is substantially less singular than its higher-dimensional counterparts: the linear stochastic convolution already has positive spatial regularity, so the cubic term can be interpreted by classical multiplication and no renormalization is needed. For our purposes, we also apply the Da Prato–Debussche decomposition $X=Y+Z$. Given an initial datum $X_{0}$, we split the solution as $X=Y+Z,$ where $Z$ is the solution of the linear stochastic equation
$dZ=-\nu\Delta^{2}Zdt+\sigma_{*}\nabla\cdot\xi(t)dt,\qquad Z(0)=0,$ (5)
and $Y$ solves the random deterministic equation
$\partial_{t}Y=-\nu\Delta^{2}Y-A\Delta(Y+Z)+\chi\,\partial_{x}^{2}\bigl((Y+Z)^{3}\bigr),\qquad Y(0)=X(0).$ (6)
In one dimension, the stochastic convolution in (5) is spatially regular enough to be treated pathwise as a function. More precisely, for every $\alpha<\frac{1}{2}$ and every $p\geq 1$, one expects
$\mathbb{E}\Bigl[\sup_{t\in[0,T]}\|Z(t)\|_{C^{\alpha}(\mathbb{T})}^{p}\Bigr]<\infty,$ (7)
and $Z\in C([0,T],C^{\alpha}(\mathbb{T})),\text{ for every }\alpha<\frac{1}{2},$ almost surely. In particular, $Z^{2}$ and $Z^{3}$ are classical pointwise products, which is the key simplification specific to the one-dimensional case.
The shifted equation for $Y$ becomes a random but classical semilinear fourth-order parabolic equation, and standard $H^{-1}$-energy estimates in $H^{-1}$, $H^{1}$, and $L^{4}$ become available. Testing (6) against $(-\partial_{x}^{2})^{-1}Y$ yields the formal coercive identity
$\frac{1}{2}\frac{d}{dt}\|Y(t)\|_{H^{-1}}^{2}+\nu\|Y(t)\|_{H^{1}}^{2}$ $=-\lambda A\,\|Y(t)\|_{L^{2}}^{2}-\lambda\chi\int_{\mathbb{T}}Y(t,x)\,(Y(t,x)+Z(t,x))^{3}\,dx.$ (8)
Using Young’s inequality,
$\int_{\mathbb{T}}Y(Y+Z)^{3}\,dx\leq-c\|Y\|_{L^{4}}^{4}+C\bigl(1+\|Z\|_{L^{\infty}}^{4}\bigr),$
so that one obtains the a priori estimate
$\sup_{t\in[0,T]}\|Y(t)\|_{H^{-1}}^{2}+\int_{0}^{T}\|Y(s)\|_{H^{1}}^{2}\,ds+\int_{0}^{T}\|Y(s)\|_{L^{4}}^{4}\,ds\leq C_{T}\Bigl(1+\|X_{0}\|_{H^{-1}}^{2}+\sup_{s\in[0,T]}\|Z(s)\|_{L^{\infty}}^{4}\Bigr).$ (9)
At the level of higher moments, one similarly expects that for every $p\geq 2$,
$\mathbb{E}\Biggl[\sup_{t\in[0,T]}\|Y(t)\|_{H^{-1}}^{2p}+\Bigl(\int_{0}^{T}\|Y(s)\|_{H^{1}}^{2}\,ds\Bigr)^{p}+\Bigl(\int_{0}^{T}\|Y(s)\|_{L^{4}}^{4}\,ds\Bigr)^{p}\Biggr]<\infty.$ (10)
We summary the well-posedness of the one-dimensional stochastic Cahn–Hilliard equation as follows.
Proposition 2.
Let $T>0$, $M\in\mathbb{R}$, $X_{0}\in H^{-1}(\mathbb{T})$ with $\int_{\mathbb{T}}X(0)dx=M$. Let $\nu>0$, $A\in\mathbb{R}$, $\chi>0$, and $\sigma_{*}\geq 0$. Then the stochastic Cahn–Hilliard equation has a unique solution $X=Y+Z$ in the class
$X\in C([0,T],H^{-1}(\mathbb{T}))\cap L^{4}((0,T)\times\mathbb{T})\quad\text{a.s.},$
with conservative total magnetization: $\int_{\mathbb{T}}X(t)dx=M,\forall t>0$.
Moreover, $X$ admits the Prato–Debussche decomposition $X=Y+Z$, where the stochastic convolution
$Z\in C([0,T],\mathcal{C}^{\alpha}(\mathbb{T}))\cap C([0,T],H^{-1}(\mathbb{T}))\quad\text{a.s.},\quad\alpha<\frac{1}{2},$
and for almost every realization of $Z$, the shifted term
$Y\in C([0,T],H^{-1}(\mathbb{T}))\cap L^{2}(0,T;H^{1}(\mathbb{T}))\cap L^{4}((0,T)\times\mathbb{T})\quad\text{a.s.}.$
For the discrete Kawasaki–Kac dynamics, the full field $X_{\gamma}$ contains a martingale part whose $H^{-1}$-Itô correction is not uniformly bounded in $\gamma$. Accordingly, the direct energy method on $X_{\gamma}$ is not the correct route. Instead, one proves convergence of the linear discrete stochastic part $Z_{\gamma}$, establishes uniform energy estimates and tightness for the deterministic remainder $Y_{\gamma}$, and finally reconstructs the limit through
$X_{\gamma}=Y_{\gamma}+Z_{\gamma}\rightarrow Y+Z=X.$
This is the conservative analogue of the shifted-equation strategy used throughout the modern analysis of weakly nonlinear stochastic evolution equations.
2.3 Dirichlet forms, entropy inequality, and the Kipnis–Varadhan bound
In this subsection we collect the canonical Gibbs measures naturally associated with the conservative Kawasaki dynamics, both on the full torus and on finite blocks, together with the corresponding Dirichlet forms and Fisher information. Standard references are Guo–Papanicolaou–Varadhan [12], Yau [22], Kipnis–Varadhan [16], and Kipnis–Landim [15].
On each invariant sector $\Sigma_{N,M}$, the canonical Gibbs measure $\mu_{N,\gamma,\beta}$ is the natural reversible reference measure.
Definition 2.1 (Global Dirichlet form on the canonical sector).
For $f:\Sigma_{N,M}\to\mathbb{R}$, define the global Dirichlet form
$D_{N,\gamma}(f):=-\bigl\langle f,\mathscr{L}_{\gamma}^{N}f\bigr\rangle_{\mu_{N,\gamma,\beta}^{M}}=N^{2}\sum_{i\in\Lambda_{N}}D_{i}^{N,\gamma},$
where the bond Dirichlet form is
$D_{i}^{N,\gamma}(f):=\frac{1}{4}\int_{\Sigma_{N,M}}c_{i}(\sigma)\,\bigl(f(\sigma^{i,i+1})-f(\sigma)\bigr)^{2}\,\mu_{N,\gamma,\beta}(d\sigma).$
The corresponding global Fisher information is $I_{N,\gamma}(f):=D_{i}^{N,\gamma}(\sqrt{f}),$ for every density $f\geq 0$ with $\int_{\Sigma_{N,M}}f\,d\mu_{N,\gamma,\beta}^{M}=1$.
Fix $\ell\geq 1$, and let $\Lambda_{\ell}:=\{-\ell,\dots,\ell\}$. The block Hamiltonian is
$H_{\gamma,\ell}^{\mathrm{blk}}(\eta):=-\frac{1}{2}\sum_{u,v\in\Lambda_{\ell}}\kappa_{\gamma}(u-v)\eta_{u}\eta_{v},\qquad\eta\in\sum_{\ell}:=\{-1,1\}^{\Lambda_{\ell}}.$
The grand canonical block Gibbs measure is
$\mu_{\gamma,\ell}^{\mathrm{gc}}(d\eta):=\frac{1}{Z_{\gamma,\ell}^{\mathrm{gc}}}\exp\!\bigl(-\beta H_{\gamma,\ell}^{\mathrm{blk}}(\eta)\bigr)\mathbf{1}_{\Sigma_{\ell}}(\eta),$
where the normalizing constant $Z_{\gamma,\ell}^{\mathrm{gc}}:=\sum_{\eta\in\Sigma_{\ell}}e^{-\beta H_{\gamma,\ell}^{\mathrm{blk}}(\eta)}.$ Although the global Kawasaki dynamics conserves only the total magnetization on the full torus, the local equilibrium appearing in the one-block replacement is naturally canonical on each block. For a prescribed block magnetization
$m\in\mathcal{M}_{\ell}:=\left\{-1,-1+\frac{2}{2\ell+1},\dots,1\right\},\qquad M_{\ell}=(2\ell+1)m,$
we define the block canonical sector
$\Sigma_{\ell,m}^{\mathrm{can}}:=\left\{\eta\in\{-1,1\}^{\Lambda_{\ell}}:\ \frac{1}{2\ell+1}\sum_{u\in\Lambda_{\ell}}\eta_{u}=m\right\}.$
Then the block canonical Gibbs measure is defined by
$\mu_{\gamma,\ell,m}^{\mathrm{can}}(d\eta):=\frac{1}{Z_{\gamma,\ell,m}^{\mathrm{can}}}\exp\!\bigl(-\beta H_{\gamma,\ell}^{\mathrm{blk}}(\eta)\bigr)\mathbf{1}_{\Sigma_{\ell,m}^{\mathrm{can}}}(\eta),$
where the normalizing constant $Z_{\gamma,\ell,m}^{\mathrm{can}}:=\sum_{\eta\in\Sigma_{\ell,m}^{\mathrm{can}}}e^{-\beta H_{\gamma,\ell}^{\mathrm{blk}}(\eta)}.$ We remark that the grand canonical block Gibbs measure can be obtain by linear combination of canonical block Gibbs measures.
Definition 2.2 (Local block Gibbs measure and local Dirichlet form).
For an internal bond $(j,j+1)$, $|j|\leq\ell-1$, define the local exchange operator
$(L_{j,j+1}^{(\ell),\gamma}g)(\eta):=c_{j,j+1}^{(\ell),\gamma}(\eta)\,\bigl(g(\eta^{j,j+1})-g(\eta)\bigr),$
and the corresponding local Dirichlet form
$D_{j,j+1}^{(\ell),\gamma}(g):=-\bigl\langle g,L_{j,j+1}^{(\ell),\gamma}g\bigr\rangle_{\mu_{\gamma,\ell}^{\mathrm{gc}}}=\frac{1}{4}\int c_{j,j+1}^{(\ell),\gamma}(\eta)\,\bigl(g(\eta^{j,j+1})-g(\eta)\bigr)^{2}\,\mu_{\gamma,\ell}^{\mathrm{gc}}(d\eta).$
Accordingly, the local Fisher information on the bond $(j,j+1)$ is $I_{j,j+1}^{(\ell),\gamma}(g):=D_{j,j+1}^{(\ell),\gamma}(\sqrt{g}).$
Lemma 5 (Uniform ellipticity of the local rates).
Under Assumption $({\bf B})$, for every fixed $\ell\geq 1$ there exist constants
$0<c_{-}(\ell,\beta)\leq c_{+}(\ell,\beta)<\infty$
such that for all $\gamma\in(0,1)$, all $|j|\leq\ell-1$, and all $\eta\in\Sigma_{\ell,m}^{\mathrm{can}}$,
$c_{-}(\ell,\beta)\leq c_{j,j+1}^{(\ell),\gamma}(\eta)\leq c_{+}(\ell,\beta).$
Similarly, there exist constants $0<\underline{c}(\beta)\leq\overline{c}(\beta)<\infty$, independent of $N$ and $\gamma$, such that
$\underline{c}(\beta)\leq c_{i}^{\gamma}(\sigma)\leq\overline{c}(\beta)$
for all $i\in\Lambda_{N}$ and all $\sigma\in\Sigma_{N,M}$.
Proof.
Since the exchange $\eta\mapsto\eta^{j,j+1}$ only modifies two spins, the corresponding energy increment
$\Delta_{j}^{(\ell)}H_{\gamma,\ell}^{\mathrm{blk}}(\eta)$
is uniformly bounded for fixed $\ell$, because $\sum_{z}\kappa_{\gamma}(z)=1$ and $|\eta_{u}|\leq 1$. The logistic form of the rate therefore yields uniform upper and lower bounds on $c_{j,j+1}^{(\ell),\gamma}(\eta)$. The same argument applies to the full-torus rates $c_{i}^{\gamma}(\sigma)$, since $|\Delta_{i}H_{\gamma}(\sigma)|$ is uniformly bounded by the $L^{1}$-normalization of the kernel. ∎
Lemma 6 (Entropy inequality).
Let $\nu$ be a probability measure, and let $\mu\ll\nu$ with relative entropy
$H(\mu\mid\nu):=\int\log\!\left(\frac{d\mu}{d\nu}\right)d\mu<\infty.$
Then for every bounded measurable random variable $F$ and every $a>0$,
$\mathbb{E}_{\mu}[F]\leq\frac{1}{a}H(\mu\mid\nu)+\frac{1}{a}\log\mathbb{E}_{\nu}[e^{aF}].$
In particular, for any random variable $Y$,
$\mathbb{E}_{\mu}[|Y|]\leq\frac{H(\mu\mid\nu)+\log 2}{a}+\frac{1}{a}\max_{\pm}\log\mathbb{E}_{\nu}[e^{\pm aY}].$
Proof.
The first inequality is the standard entropy inequality; see [15, Lemma 6.1.4] or [12, Section 2]. Applying it to $F=|Y|$ and using
$e^{a|Y|}\leq e^{aY}+e^{-aY}$
gives
$\log\mathbb{E}_{\nu}[e^{a|Y|}]\leq\log 2+\max_{\pm}\log\mathbb{E}_{\nu}[e^{\pm aY}],$
which yields the second claim. ∎
Note that the Markov generator $\mathscr{L}_{\gamma}^{N}$ is a self-adjoint non-positive operator on $L^{2}(\mu_{N,\gamma,\beta})$. We recall the standard Kipnis–Varadhan estimate for additive functionals of reversible Markov processes in the following lemma; see [16] or [15, Chapter 2, Section 1].
Lemma 7 (Kipnis–Varadhan $H^{-1}$ estimate).
Let $W:\Sigma_{N}\to\mathbb{R}$ satisfy $\int W\,d\mu_{N,\gamma,\beta}=0.$ Define the negative Sobolev norm
$\|W\|_{\mathcal{H}^{-1}_{N}}^{2}:=\sup_{f}\Bigl\{2\langle W,f\rangle_{\mu_{N,\gamma,\beta}}-\langle f,(-\mathscr{L}_{\gamma}^{N})f\rangle_{\mu_{N,\gamma,\beta}}\Bigr\}.$
Then
$\mathbb{E}_{\mu_{N,\gamma,\beta}}\left[\left(\int_{0}^{T}W(\sigma(t))\,dt\right)^{2}\right]\leq 2T\,\|W\|_{\mathcal{H}^{-1}_{N}}^{2}.$
Now we show the entropy–Kipnis–Varadhan estimate for non-equilibrium initial laws
Lemma 8.
Let $f_{N,0}\mu_{N,\gamma,\beta}$ be the initial law, and assume $H_{N}(f_{N,0}\mid\mu_{N,\gamma,\beta})\leq C_{H}N.$ Let $W:\Sigma_{N}\to\mathbb{R}$ be bounded with $\int W\,d\mu_{N,\gamma,\beta}=0.$ Under the entropy bound $H_{N}(f_{N,0}\mid\mu_{N,\gamma,\beta})\leq C_{H}N$, we have
$\mathbb{E}_{f_{N,0}}\left|\int_{0}^{T}W(\sigma(t))\,dt\right|\leq C\,\sqrt{NT}\,\|W\|_{\mathcal{H}^{-1}_{N}}.$ (11)
Proof.
Lemma 6 with $Y:=\int_{0}^{T}W(\sigma(t))\,dt,\mu=f_{N,0}\mu_{N,\gamma,\beta}$, and $\nu=\mu_{N,\gamma,\beta}$ yields that for every $a>0$,
$\mathbb{E}_{f_{N,0}}\left|\int_{0}^{T}W(\sigma(t))\,dt\right|\leq\frac{H_{N}(f_{N,0}\mid\mu_{N,\gamma,\beta})+\log 2}{a}+\frac{1}{a}\max_{\pm}\log\mathbb{E}_{\mu_{N,\gamma,\beta}}\left[\exp\!\left(\pm a\int_{0}^{T}W(\sigma(t))\,dt\right)\right].$
This gives
$\mathbb{E}_{f_{N,0}}|Y|\leq\frac{H_{N}(f_{N,0}\mid\mu_{N,\gamma,\beta})+\log 2}{a}+\frac{1}{a}\max_{\pm}\log\mathbb{E}_{\mu_{N,\gamma,\beta}}\bigl[e^{\pm aY}\bigr].$
For the exponential moment under the invariant law, one uses the standard Feynman–Kac estimate (Lemma 7) for reversible Markov processes. This yields that there exists a universal constant $C_{\mathrm{KV}}>0$ such that
$\log\mathbb{E}_{\mu_{N,\gamma,\beta}}\left[\exp\!\left(\pm a\int_{0}^{T}W(\sigma(t))\,dt\right)\right]\leq C_{\mathrm{KV}}\,a^{2}T\,\|W\|_{\mathcal{H}^{-1}_{N}}^{2}.$
Substitution gives
$\mathbb{E}_{f_{N,0}}|Y|\leq$ $\frac{H_{N}(f_{N,0}\mid\mu_{N,\gamma,\beta})+\log 2}{a}+C_{\mathrm{KV}}aT\|W\|_{\mathcal{H}^{-1}_{N}}^{2}$
$\leq$ $C\sqrt{\bigl(H_{N}(f_{N,0}\mid\mu_{N,\gamma,\beta})+1\bigr)\,T}\,\|W\|_{\mathcal{H}^{-1}_{N}}.$
Under the entropy bound $H_{N}(f_{N,0}\mid\mu_{N,\gamma,\beta})\leq C_{H}N$, this becomes estimate (11) ∎
3 Expansion for the discrete drift
In this section we give a expansion for the discrete drift for the rescaled field $X_{\gamma}$.
Since the Ising–Kac Hamiltonian
$H_{\gamma}(\sigma)=-\frac{1}{2}\sum_{x,y\in\Lambda_{N}}\kappa_{\gamma}(x-y)\sigma_{x}\sigma_{y}=-\frac{1}{2}\sum_{y\in\Lambda_{N}}h(\sigma,y)\sigma_{y},$
the exchange energy difference admits the exact representation
$\Delta_{i}H_{\gamma}(\sigma)=d_{i}\bigl(h_{\gamma}(\sigma;i)-h_{\gamma}(\sigma;i+1)\bigr)+d_{i}^{2}\,\kappa_{\gamma}(1),$
where $d_{i}:=\sigma_{i}-\sigma_{i+1}\in\{0,\pm 2\}$. Thus the nonlocal dependence of $\Delta_{i}H_{\gamma}$ enters only through the discrete gradient of the coarse-grained Kac field.
Note that the increment of $h_{\gamma}$ under a nearest-neighbour exchange:
$h_{\gamma}(\sigma^{i,i+1};k)-h_{\gamma}(\sigma;k)=d_{i}\Bigl[\kappa_{\gamma}(k-i-1)-\kappa_{\gamma}(k-i)\Bigr].$ (12)
Therefore
$(\mathscr{L}_{\gamma}^{N}h_{\gamma})(\sigma;k)=\sum_{i\in\Lambda_{N}}j_{i}(\sigma)\Bigl[\kappa_{\gamma}(k-i-1)-\kappa_{\gamma}(k-i)\Bigr].$
It is important to emphasize that the microscopic current $j_{i}$ is indexed by the bond $(i,i+1)$, not by sites $i$. Thus the left bond for each $j_{i}(\sigma)$ is not omitted; it is simply represented by the shifted index i1i-1.
Define the coarse-grained current
$\mathcal{J}_{\gamma}(t,x):=\frac{\varepsilon_{\gamma}}{\alpha_{\gamma}\delta_{\gamma}}\sum_{i\in\Lambda_{N}}\kappa_{\gamma}\!\left(\frac{x}{\varepsilon_{\gamma}}-i\right)j_{i}\!\left(\sigma\!\left(\frac{t}{\alpha_{\gamma}}\right)\right),$
where
$j_{i}(\sigma)=d_{i}\frac{1}{1+\exp{(\beta\Delta_{i}H_{\gamma}(\sigma))}}.$
Then Dynkin’s formula yields the martingale decomposition
$X_{\gamma}(t,x)=X_{\gamma}(0,x)-\int_{0}^{t}\nabla_{\varepsilon_{\gamma}}\mathcal{J}_{\gamma}(s,x)\,ds+M_{\gamma}(t,x),$
where $\nabla_{\varepsilon_{\gamma}}^{*}$ is the discrete divergence, and $M_{\gamma}$ is a pure-jump martingale. For each test function $\phi\in C^{\infty}(\mathbb{T})$, the weak drift of $\langle X_{\gamma},\phi\rangle_{\gamma}$ can be written in conservative form as
$\Bigl(\alpha_{\gamma}^{-1}\mathscr{L}_{\gamma}^{N}\langle X_{\gamma},\phi\rangle_{\gamma}\Bigr)(\sigma)=-\varepsilon_{\gamma}\sum_{i\in\Lambda_{N}}\mathcal{J}_{\gamma}(\sigma;\varepsilon_{\gamma}i)\nabla_{\varepsilon_{\gamma}}\phi(\varepsilon_{\gamma}i)=-\frac{\varepsilon_{\gamma}^{2}}{\alpha_{\gamma}\delta_{\gamma}}\sum_{i\in\Lambda_{N}}j_{i}(\sigma)\,\nabla_{\varepsilon_{\gamma}}\bigl(\mathcal{K}_{\gamma}^{\varepsilon}\phi\bigr)(\varepsilon_{\gamma}i)$
in the discrete Riemann-sum sense, where the discrete convolution operator
$(\mathcal{K}_{\gamma}^{\varepsilon}\phi)(\varepsilon_{\gamma}i):=\varepsilon_{\gamma}\sum_{j\in\Lambda_{N}}\kappa_{\gamma}(j-i)\phi(\varepsilon_{\gamma}j)=\varepsilon_{\gamma}\sum_{j\in\Lambda_{N}}\kappa_{\gamma}(i-j)\phi(\varepsilon_{\gamma}j).$
Since $X_{\gamma}(t,\varepsilon_{\gamma}i)=\delta_{\gamma}^{-1}h_{\gamma}\!\left(\sigma_{t/\alpha_{\gamma}};i\right),\varepsilon_{\gamma}i\in\mathbb{T}$, we have
$h_{\gamma}\!\left(\sigma\!\left(\frac{t}{\alpha_{\gamma}}\right);i\right)-h_{\gamma}\!\left(\sigma\!\left(\frac{t}{\alpha_{\gamma}}\right);i+1\right)=-\delta_{\gamma}\varepsilon_{\gamma}\nabla_{\varepsilon_{\gamma}}X_{\gamma}(t,\varepsilon_{\gamma}i),$
and
$\Delta_{i}H_{\gamma}=-\,d_{i}\,\delta_{\gamma}\varepsilon_{\gamma}\nabla_{\varepsilon_{\gamma}}X_{\gamma}(t,\varepsilon_{\gamma}i)+d_{i}^{2}\,\kappa_{\gamma}(1).$
Let $F(z):=\frac{1}{1+e^{z}}$, so that $j_{i}=d_{i}\,F\!\bigl(\beta\Delta_{i}H_{\gamma}\bigr).$ Set $a_{i}:=\beta d_{i}^{2}\kappa_{\gamma}(1),b_{i}:=-\beta d_{i}\,\delta_{\gamma}\varepsilon_{\gamma}\nabla_{\varepsilon_{\gamma}}X_{\gamma}(t,\varepsilon_{\gamma}i)$, then $\beta\Delta_{i}H_{\gamma}=a_{i}+b_{i}.$ Since $\kappa_{\gamma}(1)\to 0$ as $\gamma\to 0$, one has $a_{i}=O(\kappa_{\gamma}(1))$. Applying Taylor expansion to $F(a_{i}+b_{i})$ around $a_{i}$, we have
$F(a_{i}+b_{i})=F(a_{i})+F^{\prime}(a_{i})b_{i}+\frac{1}{2}F^{\prime\prime}(a_{i})b_{i}^{2}+\widetilde{R}_{i},$
with $\widetilde{R}_{i}=\frac{1}{6}F^{(3)}(a_{i}+\theta_{i}b_{i})b_{i}^{3},\ \theta_{i}\in(0,1).$ Therefore
$j_{i}=d_{i}F(a_{i})-\beta F^{\prime}(a_{i})d_{i}^{2}\,\delta_{\gamma}\varepsilon_{\gamma}\nabla_{\varepsilon_{\gamma}}X_{\gamma}+\frac{\beta^{2}}{2}F^{\prime\prime}(a_{i})d_{i}^{3}(\delta_{\gamma}\varepsilon_{\gamma}\nabla_{\varepsilon_{\gamma}}X_{\gamma})^{2}+\widetilde{R}_{i}.$
From the linear and nonlinear expansions $j_{x}$, we have decomposition $j_{i}=j_{i}^{\mathrm{lin}}+j_{i}^{\mathrm{nl}}+j_{i}^{\mathrm{rem}}.$ Here,
(i) Linear part.
Since $d_{i}^{2}\in\{0,4\}$, we have $a_{i}^{\gamma}\in\{0,4\kappa_{\gamma}(1)\}$, hence $F(a_{i}^{\gamma})$ takes only two values. Using the Taylor expansion $F(u)=\frac{1}{2}-\frac{u}{4}+O(u^{2})$ at $u=0$, one obtains
$j_{i}^{\mathrm{lin}}=\Bigl(\tfrac{1}{2}-\beta\kappa_{\gamma}(1)\Bigr)d_{i}+\widetilde{r}_{i,\gamma}^{\,\mathrm{lin}},\qquad|\widetilde{r}_{i,\gamma}^{\,\mathrm{lin}}|\lesssim\kappa_{\gamma}(1)^{2}\,|d_{i}|,$
so the leading linear current coefficient is $\tfrac{1}{2}-\beta\kappa_{\gamma}(1)$.
(ii) Nonlinear part.
Since $F^{\prime}(u)=-\frac{1}{4}+O(u)$, and $d_{i}b_{i}^{\gamma}=-d_{i}^{2}\,\delta_{\gamma}\varepsilon_{\gamma}\nabla_{\varepsilon_{\gamma}}X_{\gamma}$, we obtain the leading nonlinear contribution
$j_{i}^{\mathrm{nl}}=\frac{\beta}{4}\,d_{i}^{2}\,\delta_{\gamma}\varepsilon_{\gamma}\,\nabla_{\varepsilon_{\gamma}}X_{\gamma}(\varepsilon_{\gamma}i)+\widetilde{r}_{i,\gamma}^{\,\mathrm{nl}},\qquad|\widetilde{r}_{i,\gamma}^{\,\mathrm{nl}}|\lesssim\kappa_{\gamma}(1)\,|d_{i}|\,\delta_{\gamma}\varepsilon_{\gamma}\bigl|\nabla_{\varepsilon_{\gamma}}X_{\gamma}\bigr|.$
This is the term whose second-order Boltzmann–Gibbs closure produces the cubic drift in the limit.
(iii) Remainder.
By construction, $j_{i}^{\mathrm{rem}}=d_{i}R_{i}^{\gamma},R_{i}^{\gamma}=\frac{1}{2}F^{\prime\prime}(\theta_{i}^{\gamma})\,(b_{i}^{\gamma})^{2},$ and $F^{\prime\prime}$ is bounded on bounded sets. Hence, using $|d_{i}|\leq 2$ and $|b_{i}^{\gamma}|\lesssim\delta_{\gamma}\varepsilon_{\gamma}|\nabla_{\varepsilon_{\gamma}}X_{\gamma}|$, we have
$|j_{i}^{\mathrm{rem}}|\lesssim\delta_{\gamma}^{2}\varepsilon_{\gamma}^{2}\bigl|\nabla_{\varepsilon_{\gamma}}X_{\gamma}(\varepsilon_{\gamma}i)\bigr|^{2}.$
In particular, under the same a priori discrete $H^{1}$-type bounds used in the tightness/drift analysis, the contribution of $j_{i}^{\mathrm{rem}}$ to the weak drift is negligible.
We now lift the microscopic decomposition $j_{x}=j_{x}^{\mathrm{lin}}+j_{x}^{\mathrm{nl}}+j_{x}^{\mathrm{rem}}$ to the corresponding decomposition of the Kac coarse-grained current $\mathcal{J}_{\gamma}=\mathcal{J}_{\gamma}^{\mathrm{lin}}+\mathcal{J}_{\gamma}^{\mathrm{nl}}+\mathcal{J}_{\gamma}^{\mathrm{rem}}.$ where
$\mathcal{J}_{\gamma}^{\mathrm{lin}}(t,x)$ $:=\frac{\varepsilon_{\gamma}^{2}}{\alpha_{\gamma}\delta_{\gamma}}\sum_{i\in\Lambda_{N}}\kappa_{\gamma}\!\left(\frac{x}{\varepsilon_{\gamma}}-i\right)j_{i}^{\mathrm{lin}}\left(\sigma\!\left(\frac{t}{\alpha_{\gamma}}\right)\right)$ (13)
$\mathcal{J}_{\gamma}^{\mathrm{nl}}(t,x)$ $:=\frac{\varepsilon_{\gamma}^{2}}{\alpha_{\gamma}\delta_{\gamma}}\sum_{i\in\Lambda_{N}}\kappa_{\gamma}\!\left(\frac{x}{\varepsilon_{\gamma}}-i\right)\frac{\varepsilon_{\gamma}^{2}}{\alpha_{\gamma}\delta_{\gamma}}\,j_{i}^{\mathrm{nl}}\left(\sigma\!\left(\frac{t}{\alpha_{\gamma}}\right)\right),$ (14)
$\mathcal{J}_{\gamma}^{\mathrm{rem}}(t,x)$ $:=\frac{\varepsilon_{\gamma}^{2}}{\alpha_{\gamma}\delta_{\gamma}}\sum_{i\in\Lambda_{N}}\kappa_{\gamma}\!\left(\frac{x}{\varepsilon_{\gamma}}-i\right)j_{i}^{\mathrm{rem}}.$ (15)
Here
$|\mathcal{J}_{\gamma}^{\mathrm{rem}}(x)|\lesssim\frac{\varepsilon_{\gamma}^{4}\delta_{\gamma}}{\alpha_{\gamma}}\,\bigl|\nabla_{\varepsilon_{\gamma}}X_{\gamma}(x)\bigr|^{2}.$
In particular, (14) makes transparent why the nonlinear drift closure is governed by the second-order Boltzmann–Gibbs replacement for $d_{x}^{2}$ and by the mesoscopic replacements needed to identify $(m^{\ell})^{2}$ with $\delta_{\gamma}^{2}X_{\gamma}^{2}$: after inserting the BG closure into $\mathcal{J}_{\gamma}^{\mathrm{nl}}$ and taking the outer discrete divergence, one obtains the cubic conservative term $\partial_{x}^{2}(X^{3})$ in the scaling limit.
For each $\phi\in C^{\infty}(\mathbb{T})$, we introduce the discrete Kac-smoothed test function
$(\mathcal{K}_{\gamma}^{\varepsilon}\phi)(\varepsilon_{\gamma}i):=\varepsilon_{\gamma}\sum_{j\in\Lambda_{N}}\kappa_{\gamma}(i-j)\phi(\varepsilon_{\gamma}j).$
Then the drift term can be separated as $\alpha_{\gamma}^{-1}\mathscr{L}_{\gamma}^{N}\langle X_{\gamma},\phi\rangle_{\gamma}=\mathcal{D}_{\gamma}^{\mathrm{lin}}(\phi)+\mathcal{D}_{\gamma}^{\mathrm{nl}}(\phi)+\mathcal{D}_{\gamma}^{\mathrm{rem}}(\phi),$, where
$\mathcal{D}_{\gamma}^{\mathrm{lin}}(\phi):=\varepsilon_{\gamma}\sum_{x\in\Lambda_{\varepsilon_{\gamma}}}\phi(x)\mathcal{J}_{\gamma}^{\mathrm{lin}}(t,x)=-\Bigl(\frac{1}{2}-\beta\kappa_{\gamma}(1)\Bigr)\frac{\varepsilon_{\gamma}^{2}}{\alpha_{\gamma}\delta_{\gamma}}\sum_{x\in\Lambda_{\varepsilon_{\gamma}}}d_{x}\,\nabla_{\varepsilon_{\gamma}}(\mathcal{K}_{\gamma}^{\varepsilon}\phi)(\varepsilon_{\gamma}x)+o_{\gamma}(1),$
$\mathcal{D}_{\gamma}^{\mathrm{nl}}(\phi):=\varepsilon_{\gamma}\sum_{x\in\Lambda_{\varepsilon_{\gamma}}}\phi(x)\mathcal{J}_{\gamma}^{\mathrm{nl}}(t,x)=-\frac{\beta}{4}\frac{\varepsilon_{\gamma}^{3}}{\alpha_{\gamma}}\sum_{x\in\Lambda_{\varepsilon_{\gamma}}}d_{x}^{2}\,\nabla_{\varepsilon_{\gamma}}X_{\gamma}(\varepsilon_{\gamma}x)\,\nabla_{\varepsilon_{\gamma}}(\mathcal{K}_{\gamma}^{\varepsilon}\phi)(\varepsilon_{\gamma}x)+o_{\gamma}(1),$
$\mathcal{D}_{\gamma}^{\mathrm{rem}}(\phi):=\varepsilon_{\gamma}\sum_{x\in\Lambda_{\varepsilon_{\gamma}}}\phi(x)\mathcal{J}_{\gamma}^{\mathrm{rem}}(\varepsilon_{\gamma}x)=o_{\gamma}(1).$
Theorem 3.
For every $\phi\in C^{\infty}(\mathbb{T})$,
$\int_{0}^{t}\mathcal{D}_{\gamma}^{\mathrm{lin}}(\phi)(\sigma\big(\frac{s}{\alpha_{\gamma}}\big))\,ds=-A^{\prime}_{\gamma}\int_{0}^{t}\langle X_{\gamma}(s),\Delta_{\varepsilon_{\gamma}}\phi\rangle_{\gamma}\,ds-\nu_{\gamma}\int_{0}^{t}\langle X_{\gamma}(s),\Delta_{\varepsilon_{\gamma}}^{2}\phi\rangle_{\gamma}\,ds+o_{\gamma}(1)$
where
$\nu_{\gamma}:=\Bigl(\frac{1}{2}-\beta\kappa_{\gamma}(1)\Bigr)\frac{\mathfrak{m}_{2,\gamma}}{2}\frac{\varepsilon_{\gamma}^{4}}{\alpha_{\gamma}\gamma^{2}}\to\frac{\lambda}{4}\mathfrak{m}_{2}=\nu,\quad\text{and}\quad A^{\prime}_{\gamma}:=\Bigl(\frac{1}{2}-\beta\kappa_{\gamma}(1)\Bigr)\frac{\varepsilon_{\gamma}^{2}}{\alpha_{\gamma}}\to\frac{1}{2}\frac{\varepsilon_{\gamma}^{2}}{\alpha_{\gamma}}:=A^{\prime}.$
Proof.
Recall that the linear part is defined as
$\mathcal{D}_{\gamma}^{\mathrm{lin}}(\phi):=-\Bigl(\frac{1}{2}-\beta\kappa_{\gamma}(1)\Bigr)\frac{\varepsilon_{\gamma}^{2}}{\alpha_{\gamma}\delta_{\gamma}}\sum_{x\in\Lambda_{\varepsilon_{\gamma}}}d_{x}\,\nabla_{\varepsilon_{\gamma}}(\mathcal{K}_{\gamma}^{\varepsilon}\phi)(x)+o_{\gamma}(1)$ (16)
Note that $d_{x}=\sigma_{x}-\sigma_{x+1}$ is the conservative gradient of the spin field, and $X_{\gamma}$ is defined by the coarse-grained and rescale of the spin field. Using the discrete integration by parts, the linear part $\mathcal{D}_{\gamma}^{\mathrm{lin}}(\phi)$ can be rewritten as
$\mathcal{D}_{\gamma}^{\mathrm{lin}}(\phi)=-\Bigl(\frac{1}{2}-\beta\kappa_{\gamma}(1)\Bigr)\frac{\varepsilon_{\gamma}^{2}}{\alpha_{\gamma}}\langle X_{\gamma},\Delta_{\varepsilon_{\gamma}}(\mathcal{K}_{\gamma}^{\varepsilon}\phi)\rangle_{\gamma}+o_{\gamma}(1).$ (17)
Here the factor $\delta_{\gamma}^{-1}$ is exactly absorbed into $X_{\gamma}$.
We now expand the operator $\mathcal{K}_{\gamma}^{\varepsilon}$ on smooth functions. Since $\kappa_{\gamma}$ is even and has finite second and fourth moments, a standard discrete moment expansion gives
$\mathcal{K}_{\gamma}^{\varepsilon}\phi=\phi+\frac{\mathfrak{m}_{2,\gamma}}{2}\Bigl(\frac{\varepsilon_{\gamma}}{\gamma}\Bigr)^{2}\Delta_{\varepsilon_{\gamma}}\phi+r_{\gamma}(\phi),$ (18)
where
$\mathfrak{m}_{2,\gamma}:=\gamma^{2}\sum_{z\in\mathbb{Z}}z^{2}\,\kappa_{\gamma}(z)=\gamma\sum_{z\in\mathbb{Z}}(\gamma z)^{2}\,\mathfrak{K}(\gamma z),$
and the remainder satisfies $\|r_{\gamma}(\phi)\|_{\infty}\leq C_{\phi}\Bigl(\frac{\varepsilon_{\gamma}}{\gamma}\Bigr)^{4}.$ Applying $\Delta_{\varepsilon_{\gamma}}$ to (18), we obtain
$\Delta_{\varepsilon_{\gamma}}(\mathcal{K}_{\gamma}^{\varepsilon}\phi)=\Delta_{\varepsilon_{\gamma}}\phi+\frac{\mathfrak{m}_{2}}{2}\Bigl(\frac{\varepsilon_{\gamma}}{\gamma}\Bigr)^{2}\Delta_{\varepsilon_{\gamma}}^{2}\phi+\Delta_{\varepsilon_{\gamma}}r_{\gamma}(\phi).$ (19)
Since $\phi$ is smooth, the discrete Laplacian of the remainder satisfies the same order bound:
$\|\Delta_{\varepsilon_{\gamma}}r_{\gamma}(\phi)\|_{\infty}\leq C_{\phi}\Bigl(\frac{\varepsilon_{\gamma}}{\gamma}\Bigr)^{4}.$ (20)
Substituting (19) into (17) yields
$\mathcal{D}_{\gamma}^{\mathrm{lin}}(\phi)=-\Bigl(\frac{1}{2}-\beta\kappa_{\gamma}(1)\Bigr)\frac{\varepsilon_{\gamma}^{2}}{\alpha_{\gamma}}\langle X_{\gamma},\Delta_{\varepsilon_{\gamma}}\phi\rangle_{\gamma}-\Bigl(\frac{1}{2}-\beta\kappa_{\gamma}(1)\Bigr)\frac{\mathfrak{m}_{2,\gamma}}{2}\frac{\varepsilon_{\gamma}^{4}}{\alpha_{\gamma}\gamma^{2}}\langle X_{\gamma},\Delta_{\varepsilon_{\gamma}}^{2}\phi\rangle_{\gamma}-\frac{\varepsilon_{\gamma}^{2}}{\alpha_{\gamma}}\langle X_{\gamma},\Delta_{\varepsilon_{\gamma}}r_{\gamma}(\phi)\rangle_{\gamma}.$
Here, the first two terms in above identity are $A^{\prime}_{\gamma}\langle X_{\gamma},\Delta_{\varepsilon_{\gamma}}\phi\rangle_{\gamma}-\nu_{\gamma}\langle X_{\gamma},\Delta_{\varepsilon_{\gamma}}^{2}\phi\rangle_{\gamma}.$. It remains to prove that the last term is negligible.
By the weak $L^{2}$-bound on $X_{\gamma}$ and the uniform bound (20), we have
$\mathbb{E}\Bigg[\Bigg|c_{\gamma}^{\mathrm{lin}}\frac{\varepsilon_{\gamma}^{2}}{\alpha_{\gamma}}\int_{0}^{t}\langle X_{\gamma}(s),\Delta_{\varepsilon_{\gamma}}r_{\gamma}(\phi)\rangle_{\gamma}\,ds\Bigg|^{2}\Bigg]\leq C_{t}\Bigl(c_{\gamma}^{\mathrm{lin}}\frac{\varepsilon_{\gamma}^{2}}{\alpha_{\gamma}}\Bigr)^{2}\|\Delta_{\varepsilon_{\gamma}}r_{\gamma}(\phi)\|_{\infty}^{2}\leq C_{t,\phi}\Bigl(\frac{\varepsilon_{\gamma}^{2}}{\alpha_{\gamma}}\Bigr)^{2}\Bigl(\frac{\varepsilon_{\gamma}}{\gamma}\Bigr)^{8}.$
Using $\frac{\varepsilon_{\gamma}^{2}}{\alpha_{\gamma}}=\frac{A^{\prime}_{\gamma}}{\frac{1}{2}-\beta\kappa_{\gamma}(1)}$, and the boundedness of $\frac{\varepsilon_{\gamma}^{2}}{\alpha_{\gamma}}$, we have $\lim_{\gamma\to 0}\frac{\varepsilon_{\gamma}^{6}}{\alpha_{\gamma}\gamma^{4}}=0.$
Hence as $\gamma\to 0$,
$c_{\gamma}^{\mathrm{lin}}\frac{\varepsilon_{\gamma}^{2}}{\alpha_{\gamma}}\int_{0}^{t}\langle X_{\gamma}(s),\Delta_{\varepsilon_{\gamma}}r_{\gamma}(\phi)\rangle_{\gamma}\,ds\longrightarrow 0\qquad\text{in }L^{2}.$
Therefore
$\int_{0}^{t}\mathcal{D}_{\gamma}^{\mathrm{lin}}(\phi)(\sigma\big(\frac{t}{\alpha_{\gamma}}\big))\,ds=-A^{\prime}_{\gamma}\int_{0}^{t}\langle X_{\gamma}(s),\Delta_{\varepsilon_{\gamma}}\phi\rangle_{\gamma}\,ds-\nu_{\gamma}\int_{0}^{t}\langle X_{\gamma}(s),\Delta_{\varepsilon_{\gamma}}^{2}\phi\rangle_{\gamma}\,ds+o_{\gamma}(1)$
in $L^{2}$, as claimed. ∎
By Theorem 3, the rescaled field $X_{\gamma}$ satisfies the following discrete SPDE on lattice $\Lambda_{N}$:
$dX_{\gamma}=\Bigl[-\nu_{\gamma}\Delta_{\varepsilon_{\gamma}}^{2}X_{\gamma}-A^{\prime}_{\gamma}\Delta_{\varepsilon_{\gamma}}X_{\gamma}+\mathcal{N}(X_{\gamma})+\mathcal{R}(X_{\gamma})\Bigr]dt+dM_{\gamma}.$
where $\mathcal{N}(X_{\gamma}):=\nabla_{\varepsilon_{\gamma}}\mathcal{J}_{\gamma}^{\mathrm{nl}}$, and the remainder term $\mathcal{R}(X_{\gamma}):=\nabla_{\varepsilon_{\gamma}}\mathcal{J}_{\gamma}^{\mathrm{rem}}\to 0$ as $\gamma\to 0$.
In order to obtain the uniform moment bound and the uniform drift bound needed in the tightness argument, it is more convenient to work with the Da Prato–Debussche type decomposition of the rescaled field rather than to apply Itô’s formula directly to the full process $X_{\gamma}$.
By using the discrete semigroup $S_{\gamma}:=e^{-\nu_{\gamma}t\Delta^{2}_{\varepsilon_{\gamma}}}$, we define the discrete stochastic convolution $Z_{\gamma}:=\int^{t}_{0}S_{\gamma}(t-s)M_{\gamma}(t)$, so that
$dZ_{\gamma}=-\nu_{\gamma}\Delta_{\varepsilon_{\gamma}}^{2}Z_{\gamma}dt+dM_{\gamma},\qquad Z_{\gamma}(0)=0,$ (21)
Then we introduce the Da Prato–Debussche type decompose $X_{\gamma}=Y_{\gamma}+Z_{\gamma}.$ Here, the shifted field $Y_{\gamma}$ solves the random deterministic equation
$\partial_{t}Y_{\gamma}=-\nu_{\gamma}\Delta_{\varepsilon_{\gamma}}^{2}Y_{\gamma}-A^{\prime}_{\gamma}Y_{\gamma}+\mathcal{N}_{\gamma}(X_{\gamma})+\mathcal{R}_{\gamma}(X_{\gamma}),\qquad Y_{\gamma}(0)=X_{\gamma}(0).$ (22)
4 The Second-order Boltzmann–Gibbs Principle
Throughout this section, all limits are taken along a sequence $\gamma\downarrow 0,N=N_{\gamma}\to\infty.$ We denote by $\mu_{N,\gamma,\beta}$ the corresponding reversible Gibbs measure of the one-dimensional Ising–Kac–Kawasaki dynamics $\sigma_{t}$.
We assume that the initial law is $f_{N,0}\mu_{N,\gamma,\beta}$, where
$H_{N}(f_{N,0}\mid\mu_{N,\gamma,\beta})\leq CN$
for some constant $C>0$, and that the initial empirical magnetization field converges to a macroscopic profile $m_{0}$.
Fix $r\geq 1$, and let $\Psi:\{-1,1\}^{\Lambda_{r}}\to\mathbb{R}$ be a bounded local observable. For $\ell>r$, define the centered block
$\Lambda_{\ell}:=\{-\ell,\dots,\ell\},\qquad\Lambda_{\ell}(x):=x+\Lambda_{\ell},$
and the corresponding block average
$m_{x}^{\ell}(\sigma):=\frac{1}{2\ell+1}\sum_{y\in\Lambda_{\ell}(x)}\sigma_{y}.$
With $L\gg\ell$, we also define the mesoscopic block average in two block estimates
$m_{x}^{L}(\sigma):=\frac{1}{2L+1}\sum_{y\in\Lambda_{L}(x)}\sigma_{y},\qquad\Lambda_{L}(x):=x+\{-L,\dots,L\}.$
For each bond $z=(z,z+1)\in\Lambda_{N}$, define $A_{z}(\sigma):=\sigma(z+1)-\sigma(z),$ and let
$D_{z}^{N,\gamma}(f):=\frac{1}{4}\int c_{z}(\sigma)\,\bigl(f(\sigma^{z,z+1})-f(\sigma)\bigr)^{2}\,\mu_{N,\gamma,\beta}(d\sigma)$
be the local Dirichlet form associated with the Kawasaki exchange on the bond $(z,z+1)$.
In this section, we also assume that following assumption hold.
Assumption 4 (Block scales).
There exist mesoscopic block sizes $\ell=\ell_{\gamma},L:=L_{\gamma}\in\mathbb{N}$ such that
$\ell=\ell(\gamma)\to\infty,\quad L=L(\gamma)\to\infty,\quad\frac{\ell}{L}\to 0,\quad\frac{L}{N}\to 0,\quad\gamma\ell\to 0,\quad\gamma^{2}L^{3}\to 0.$
Whenever the weighted block-to-Kac replacement is used, we further assume
$\sup_{\gamma\downarrow 0}\frac{\varepsilon_{\gamma}^{2}\gamma^{1/2}}{\alpha_{\gamma}\delta_{\gamma}}<\infty.$
4.1 One-block replacement
On each reference block $\Lambda_{\ell}$, we consider the bounded local observable $\Psi:\{-1,1\}^{\mathbb{Z}}\to\mathbb{R}$ with
$\mathrm{supp}(\Psi)\subset\Lambda_{r}:=\{-r,\dots,r\}$
for some $r\geq 1$. Using the local canonical Gibbs measure $\mu_{\gamma,\ell,m}^{\mathrm{can}}(d\eta)$, for every $\ell\geq r$, we define
$\Phi_{\Psi,\ell}^{\gamma}(m):=\int\Psi(\eta)\,\mu_{\gamma,\ell,m}^{\mathrm{can}}(d\eta),\quad m\in\mathcal{M}_{\ell}.$
We have the following one-block estimate.
Theorem 5 (One-block estimate).
For every smooth test function $J\in C^{\infty}(\mathbb{T})$, as $\ell=\ell(\gamma)\to\infty,\gamma\ell(\gamma)\to 0$,
$\lim_{\ell_{\gamma}\to\infty}\limsup_{\gamma\ell_{\gamma}\to 0}\lim_{N\to\infty}\mathbb{E}_{f_{N,0}}\left|\int_{0}^{T}\frac{1}{N}\sum_{x\in\Lambda_{N}}J\!\left(\frac{x}{N}\right)\Bigl(\tau_{x}\Psi(\sigma(t))-\Phi_{\Psi,\ell}^{\gamma}(m_{x}^{\ell}(\sigma(t)))\Bigr)\,dt\right|=0.$
Proof.
Let $f_{N,t}$ be the density of the law of $\sigma(t)$ with respect to the canonical Gibbs measure $\mu_{N,\gamma,\beta}$. By reversibility and entropy dissipation,
$\frac{d}{dt}H_{N}(f_{N,t}\mid\mu_{N,\gamma,\beta})=-4N^{2}\sum_{x\in\Lambda_{N}}I_{x,x+1}^{N,\gamma}(f_{N,t}),$
where
$I_{x,x+1}^{N,\gamma}(f):=\frac{1}{4}\int c_{x,x+1}^{\gamma}(\sigma)\Bigl(\sqrt{f(\sigma^{x,x+1})}-\sqrt{f(\sigma)}\Bigr)^{2}\,\mu_{N,\gamma,\beta}(d\sigma).$
Hence
$H_{N}(f_{N,t}\mid\mu_{N,\gamma,\beta})\leq CN,\qquad\int_{0}^{T}\sum_{x\in\Lambda_{N}}I_{x,x+1}^{N,\gamma}(f_{N,t})\,dt\leq\frac{C}{N}.$
Define the time-averaged density $\bar{f}_{N}:=\frac{1}{T}\int_{0}^{T}f_{N,t}\,dt.$ By convexity of entropy and Fisher information,
$H_{N}(\bar{f}_{N}\mid\mu_{N,\gamma,\beta})\leq CN,\qquad\sum_{x\in\Lambda_{N}}I_{x,x+1}^{N,\gamma}(\bar{f}_{N})\leq\frac{C}{N}.$ (23)
For fixed $\ell$, define the local block projection $\pi_{x,\ell}:\Sigma_{N}\to\{-1,1\}^{\Lambda_{\ell}},\pi_{x,\ell}(\sigma):=(\sigma_{x+y})_{y\in\Lambda_{\ell}},$ and the spatially averaged local law
$Q_{N}^{\ell}(d\theta,d\eta):=d\theta\,\beta_{\theta}^{N,\ell}(d\eta)=\frac{1}{2N+1}\sum_{x\in\Lambda_{N}}\delta_{x/N}(d\theta)\,(\pi_{x,\ell})_{*}(\bar{f}_{N}\mu_{N,\gamma,\beta})(d\eta),\quad(\theta,\eta)\in\mathbb{T}\times\{-1,1\}^{\Lambda_{\ell}}.$
Passing to a subsequence if necessary, we may assume
$Q_{N}^{\ell}\rightarrow Q^{\ell}\ \text{as}\ N\to\infty,\quad Q^{\ell}(d\theta,d\eta)=d\theta\,\beta_{\theta}^{\ell}(d\eta),$
where $\beta_{\theta}^{\ell}$ is a probability kernel on $\{-1,1\}^{\Lambda_{\ell}}$.
Now we claim that for every fixed $\ell$,
$I_{j,j+1}^{(\ell),\gamma}\!\left(g_{x,\ell}\right)\,d\theta=\sum_{|j|\leq\ell-1}\int_{\mathbb{T}}I_{j,j+1}^{(\ell),\gamma}\!\left(\frac{d\beta_{\theta}^{N,\ell}}{d\mu_{\gamma,\ell}^{\mathrm{gc}}}\right)\,d\theta\leq C_{\ell}\sum_{x\in\Lambda_{N}}I_{x,x+1}^{N,\gamma}(\bar{f}_{N}),$ (24)
where $\mu_{\gamma,\ell}^{\mathrm{gc}}(d\eta)$ is the grand canonical Gibbs reference measure on the block $\Lambda_{\ell}$, and $g_{x,\ell}$ is the density $d\beta_{\theta}^{N,\ell}/d\mu_{\gamma,\ell}^{\mathrm{gc}}$ on the block $\Lambda_{\ell}$. To see this claim, we need show the monotonicity of the projected Fisher information:
$\widehat{I}_{x,j}(g_{x,\ell}):=\frac{1}{4}\int\widehat{c}_{x,j}(\eta)\Bigl(\sqrt{g_{x,\ell}(\eta^{j,j+1})}-\sqrt{g_{x,\ell}(\eta)}\Bigr)^{2}\,\widehat{\mu}_{x,\ell}(d\eta).$
We first show that projection does not increase the local Fisher information for the exact block marginal, i.e.
$\widehat{I}_{x,j}(g_{x,\ell})\leq I_{x+j,x+j+1}^{N,\gamma}(\bar{f}_{N}).$
To prove this, write a global configuration as $\sigma=(\eta,\xi)$, where $\eta$ denotes the block variables and $\xi$ the exterior variables. Then $g_{x,\ell}(\eta)$ is obtained by conditional averaging of $\bar{f}_{N}(\eta,\xi)$ over $\xi$, and the exchange $\eta\mapsto\eta^{j,j+1}$ corresponds exactly to the global exchange on the bond $(x+j,x+j+1)$.
For each fixed $\eta$, we have
$\Bigl(\sqrt{g_{x,\ell}(\eta^{j,j+1})}-\sqrt{g_{x,\ell}(\eta)}\Bigr)^{2}\leq\int\Bigl(\sqrt{\bar{f}_{N}(\eta^{j,j+1},\xi)}-\sqrt{\bar{f}_{N}(\eta,\xi)}\Bigr)^{2}\,\mu_{N,\gamma,\beta}(d\xi\mid\eta).$
Multiplying by $\widehat{c}_{x,j}(\eta)$, integrating over $\eta$ with respect to $\widehat{\mu}_{x,\ell}$, and using the very definition of $\widehat{c}_{x,j}$ as a conditional expectation of the global rate, we obtain
$\widehat{I}_{x,j}(g_{x,\ell})\leq\frac{1}{4}\int c_{x+j,x+j+1}^{\gamma}(\sigma)\Bigl(\sqrt{\bar{f}_{N}(\sigma^{x+j,x+j+1})}-\sqrt{\bar{f}_{N}(\sigma)}\Bigr)^{2}\,\mu_{N,\gamma,\beta}(d\sigma)=I_{x+j,x+j+1}^{N,\gamma}(\bar{f}_{N}).$
Since $\ell$ is fixed, the block state space $\{-1,1\}^{\Lambda_{\ell}}$ is finite. Moreover, by the boundedness of the Kac interaction on a finite block, the exact block marginal $\widehat{\mu}_{x,\ell}$ and the reference grand canonical block Gibbs measure $\mu_{\gamma,\ell}^{\mathrm{gc}}$ are uniformly equivalent. Moreover, the exact projected rates $\widehat{c}_{x,j}$ and the standard block rates $c_{j,j+1}^{(\ell),\gamma}$ defining $I_{j,j+1}^{(\ell),\gamma}$ are uniformly comparable. Thus, there exists $C_{\ell}$ such that
$I_{j,j+1}^{(\ell),\gamma}\!\left(\frac{d\beta_{x/N}^{N,\ell}}{d\mu_{\gamma,\ell}^{\mathrm{gc}}}\right)\leq C_{\ell}\,\widehat{I}_{x,j}(g_{x,\ell})\leq C_{\ell}I_{x+j,x+j+1}^{N,\gamma}(\bar{f}_{N}).$
Averaging over $x\in\Lambda_{N}$ and summing over $|j|\leq\ell-1$, we obtain
$\sum_{|j|\leq\ell-1}\int_{\mathbb{T}}I_{j,j+1}^{(\ell),\gamma}\!\left(\frac{d\beta_{\theta}^{N,\ell}}{d\mu_{\gamma,\ell}^{\mathrm{gc}}}\right)\,d\theta\leq C_{\ell}\sum_{|j|\leq\ell-1}\frac{1}{N}\sum_{x\in\Lambda_{N}}I_{x+j,x+j+1}^{N,\gamma}(\bar{f}_{N}).$
Now fix a global bond $(x+j,x+j+1)$. Hence each global bond is counted at most $2\ell$ times in the double sum, and therefore
$\sum_{|j|\leq\ell-1}\frac{1}{N}\sum_{x\in\Lambda_{N}}I_{x+j,x+j+1}^{N,\gamma}(\bar{f}_{N})\leq C_{\ell}^{(4)}\sum_{b\in\Lambda_{N}}I_{b,b+1}^{N,\gamma}(\bar{f}_{N}).$
Combining the last two displays yields (24).
Then combining with the estimate (23), we obtain
$\sum_{|j|\leq\ell-1}\int_{\mathbb{T}}I_{j,j+1}^{(\ell),\gamma}\!\left(\frac{d\beta_{\theta}^{N,\ell}}{d\mu_{\gamma,\ell}^{\mathrm{gc}}}\right)\,d\theta\leq\frac{C_{\ell}}{N}\to 0.$
By lower semicontinuity, for almost every $\theta$,
$I_{j,j+1}^{(\ell),\gamma}\!\left(\frac{d\beta_{\theta}^{\ell}}{d\mu_{\gamma,\ell}^{\mathrm{gc}}}\right)=0,\qquad|j|\leq\ell-1.$
Hence $d\beta_{\theta}^{\ell}/d\mu_{\gamma,\ell}^{\mathrm{gc}}$ is invariant under all internal nearest-neighbour exchanges. Since these exchanges preserve the block magnetization and the exchange graph is connected on each fixed-magnetization sector, $\beta_{\theta}^{\ell}$ must be a convex combination of the canonical block Gibbs measures:
$\beta_{\theta}^{\ell}=\sum_{m\in\mathcal{M}_{\ell}}\vartheta_{\theta}^{\ell}(m)\,\mu_{\gamma,\ell,m}^{\mathrm{can}},\qquad\mathcal{M}_{\ell}=\left\{-\!1,-\!1+\frac{2}{2\ell+1},\dots,1\right\}.$
Note that $\Psi$ is supported in $\Lambda_{r}\subset\Lambda_{\ell}$ and $m^{\ell}(\eta):=\frac{1}{2\ell+1}\sum_{u\in\Lambda_{\ell}}\eta_{u}$. It follows that
$\int_{\sum_{\ell}}\Bigl(\Psi(\eta)-\Phi_{\Psi,\ell}^{\gamma}(m^{\ell}(\eta))\Bigr)\,\beta_{\theta}^{\ell}(d\eta)=0.$
Since $Q_{N}^{\ell}\rightarrow Q^{\ell}$ as $N\to\infty$, we have
$\lim_{N\to\infty}\int J(\theta)\Bigl(\Psi(\eta)-\Phi_{\Psi,\ell}^{\gamma}(m^{\ell}(\eta))\Bigr)dQ_{N}^{\ell}=\int_{\mathbb{T}}J(\theta)\left[\int\Bigl(\Psi(\eta)-\Phi_{\Psi,\ell}^{\gamma}(m^{\ell}(\eta))\Bigr)\,\beta_{\theta}^{\ell}(d\eta)\right]d\theta=0.$
By the definition of $Q_{N}^{\ell}$, this means that for every fixed $\ell$,
$\lim_{N\to\infty}\mathbb{E}_{f_{N,0}}\left[\int_{0}^{T}\frac{1}{N}\sum_{x\in\Lambda_{N}}J\!\left(\frac{x}{N}\right)\Bigl(\tau_{x}\Psi(\sigma(t))-\Phi_{\Psi,\ell}^{\gamma}(m_{x}^{\ell}(\sigma(t)))\Bigr)\,dt\right]=0.$
Since $\Psi$ and $\Phi_{\Psi,\ell}^{\gamma}$ are uniformly bounded, the random variables above are uniformly bounded in $L^{\infty}$, so convergence of expectations implies convergence of the expectation of the absolute value. Finally, letting $\ell\to\infty$ with $\gamma\ell\to 0$ gives the claim. ∎
4.2 Two-block replacement for the conserved field
We first show the follow local $H^{-1}$ bound for bond differences.
Lemma 9.
There exists a constant $C>0$, depending only on $c_{-}$ and $c_{+}$, such that for every $z\in\Lambda_{N}$ and every bounded function $f:\Sigma_{N}\to\mathbb{R}$,
$\bigl|\langle A_{z},f\rangle_{\mu_{N,\gamma,\beta}}\bigr|\leq C\,\sqrt{D_{z}^{N,\gamma}(f)}.$
Consequently, for every family of coefficients $(b_{z})_{z\in\Lambda_{N}}$, the function $W(\sigma):=\sum_{z\in\Lambda_{N}}b_{z}\,A_{z}(\sigma)$ satisfies
$\|W\|_{\mathcal{H}^{-1}_{N}}^{2}\leq\frac{C}{N^{2}}\sum_{z\in\Lambda_{N}}|b_{z}|^{2},$ (25)
where
$\|W\|_{\mathcal{H}^{-1}_{N}}^{2}:=\sup_{f}\Bigl\{2\langle W,f\rangle_{\mu_{N,\gamma,\beta}}-N^{2}D_{N}^{N,\gamma}(f)\Bigr\},\quad D_{N}^{N,\gamma}(f):=\sum_{z\in\Lambda_{N}}D_{z}^{N,\gamma}(f).$
Proof.
We first prove the local estimate for a single bond. For a fixed bond $z=(z,z+1)$, let
$(L_{z}^{N,\gamma}g)(\sigma):=c_{z}(\sigma)\bigl(g(\sigma^{z,z+1})-g(\sigma)\bigr)$
be the local exchange generator on that bond. We claim that there exists a bounded function $h_{z}:\Sigma_{N}\to\mathbb{R}$ such that $L_{z}^{N,\gamma}h_{z}=A_{z}.$ If $\sigma^{z,z+1}=\sigma$, then $\sigma(z)=\sigma(z+1)$, hence $A_{z}(\sigma)=\sigma(z+1)-\sigma(z)=0,$ and there is no constraint on $h_{z}$ on this orbit. If $\sigma^{z,z+1}\neq\sigma$, then the orbit consists of the pair $\{\sigma,\sigma^{z,z+1}\},$ and $A_{z}(\sigma^{z,z+1})=-A_{z}(\sigma).$ Choose arbitrarily one representative $\sigma$ from the orbit and set $h_{z}(\sigma)=0$. Then define
$h_{z}(\sigma^{z,z+1}):=\frac{A_{z}(\sigma)}{c_{z}(\sigma)}.$
By construction, we have
$(L_{z}^{N,\gamma}h_{z})(\sigma)=c_{z}(\sigma)\bigl(h_{z}(\sigma^{z,z+1})-h_{z}(\sigma)\bigr)=A_{z}(\sigma),$
For the swapped configuration $\sigma^{z,z+1}$, using detailed balance and the antisymmetry of $A_{z}$, the same identity holds automatically. Hence $L_{z}^{N,\gamma}h_{z}=A_{z}$ on the whole state space.
Moreover, since $|A_{z}(\sigma)|\leq 2$, by Lemma 5 we have
$|h_{z}(\sigma^{z,z+1})-h_{z}(\sigma)|=\left|\frac{A_{z}(\sigma)}{c_{z}(\sigma)}\right|\leq\frac{2}{c_{-}}.$ (26)
Since $\mu_{N,\gamma,\beta}$ is reversible for the Kawasaki dynamics, for every pair of bounded functions $g,f$,
$\langle L_{z}^{N,\gamma}g,f\rangle_{\mu_{N,\gamma,\beta}}=-\frac{1}{2}\int c_{z}(\sigma)\,\bigl(g(\sigma^{z,z+1})-g(\sigma)\bigr)\bigl(f(\sigma^{z,z+1})-f(\sigma)\bigr)\,\mu_{N,\gamma,\beta}(d\sigma).$
Applying this with $g=h_{z}$ and using $L_{z}^{N,\gamma}h_{z}=A_{z}$, we obtain
$\langle A_{z},f\rangle_{\mu_{N,\gamma,\beta}}=-\frac{1}{2}\int c_{z}(\sigma)\,\bigl(h_{z}(\sigma^{z,z+1})-h_{z}(\sigma)\bigr)\bigl(f(\sigma^{z,z+1})-f(\sigma)\bigr)\,\mu_{N,\gamma,\beta}(d\sigma).$
By Cauchy–Schwarz inequality and the bound (26), we have
$\bigl|\langle A_{z},f\rangle_{\mu_{N,\gamma,\beta}}\bigr|$
$\leq$ $\left(\frac{1}{2}\int c_{z}(\sigma)\,\bigl(h_{z}(\sigma^{z,z+1})-h_{z}(\sigma)\bigr)^{2}\,d\mu_{N,\gamma,\beta}\right)^{1/2}\left(\frac{1}{2}\int c_{z}(\sigma)\,\bigl(f(\sigma^{z,z+1})-f(\sigma)\bigr)^{2}\,d\mu_{N,\gamma,\beta}\right)^{1/2}$
$\leq$ $C\sqrt{D_{z}^{N,\gamma}(f)},$
which proves the first claim. Then, by the local bound just proved, we have
$|\langle W,f\rangle_{\mu_{N,\gamma,\beta}}|\leq\sum_{z\in\Lambda_{N}}|b_{z}|\,|\langle A_{z},f\rangle_{\mu_{N,\gamma,\beta}}|\leq C\sum_{z\in\Lambda_{N}}|b_{z}|\,\sqrt{D_{z}^{N,\gamma}(f)}\leq N^{2}D_{N}^{N,\gamma}(f)+\frac{C}{N^{2}}\sum_{z\in\Lambda_{N}}|b_{z}|^{2}.$
Using the variational definition of the $\mathcal{H}^{-1}_{N}$-norm, we conclude (25). This proves the lemma. ∎
Theorem 6 (Two-block estimate).
For every smooth test function $J\in C^{\infty}(\mathbb{T})$,
$\lim_{\gamma\downarrow 0}\mathbb{E}_{f_{N,0}}\left|\int_{0}^{T}\frac{1}{N}\sum_{x\in\Lambda_{N}}J\!\left(\frac{x}{N}\right)\Bigl(m_{x}^{\ell_{\gamma}}(\sigma(t))-m_{x}^{L_{\gamma}}(\sigma(t))\Bigr)\,dt\right|=0.$
Proof.
We split the argument into two steps.
Step 1: from $m_{x}^{\ell}$ to the sliding average of $\ell$-blocks. Define
$\widetilde{m}_{x}^{\ell,L}(\sigma):=\frac{1}{2L+1}\sum_{|y|\leq L}m_{x+y}^{\ell}(\sigma).$
Then
$m_{x}^{\ell}-\widetilde{m}_{x}^{\ell,L}=\frac{1}{2L+1}\sum_{|y|\leq L}\bigl(m_{x}^{\ell}-m_{x+y}^{\ell}\bigr).$
Hence it suffices to prove
$\lim_{\gamma\downarrow 0}\mathbb{E}_{f_{N,0}}\left|\int_{0}^{T}\frac{1}{N}\sum_{x\in\Lambda_{N}}J\!\left(\frac{x}{N}\right)\Bigl(m_{x}^{\ell}(\sigma(t))-\widetilde{m}_{x}^{\ell,L}(\sigma(t))\Bigr)\,dt\right|=0.$
Fix $y>0$. By definition,
$m_{x+y}^{\ell}-m_{x}^{\ell}=\frac{1}{2\ell+1}\sum_{u=-\ell}^{\ell}\bigl(\sigma_{x+y+u}-\sigma_{x+u}\bigr).$
Note that the exchange between $x+u$ and $x+y+u$ can be realized telescopically by the exchanges between neighboring sites from $x+u$ to $x+y+u$. Using the telescoping identity
$\sigma_{x+y+u}-\sigma_{x+u}=\sum_{k=0}^{y-1}\bigl(\sigma_{x+u+k+1}-\sigma_{x+u+k}\bigr),$
we obtain
$m_{x+y}^{\ell}-m_{x}^{\ell}=\frac{1}{2\ell+1}\sum_{u=-\ell}^{\ell}\sum_{k=0}^{y-1}\bigl(\sigma_{x+u+k+1}-\sigma_{x+u+k}\bigr).$
The same formula holds for $y<0$ after reversing the path.
Thus
$\frac{1}{N}\sum_{x\in\Lambda_{N}}J\!\left(\frac{x}{N}\right)\bigl(m_{x+y}^{\ell}-m_{x}^{\ell}\bigr)=\sum_{z\in\Lambda_{N}}b_{z}^{N,\ell,y}\bigl(\sigma(z+1)-\sigma(z)\bigr),$
for deterministic coefficients $b_{z}^{N,\ell,y}$. Averaging over $|y|\leq L$, we obtain
$\frac{1}{N}\sum_{x\in\Lambda_{N}}J\!\left(\frac{x}{N}\right)\bigl(m_{x}^{\ell}-\widetilde{m}_{x}^{\ell,L}\bigr)=\sum_{z\in\Lambda_{N}}b_{z}^{N,\ell,L}\bigl(\sigma(z+1)-\sigma(z)\bigr),$
with coefficients satisfying the counting bound $\sum_{z\in\Lambda_{N}}|b_{z}^{N,\ell,L}|^{2}\leq C_{J}\,\frac{L}{N}.$ Let
$W_{N,\ell,L}(\sigma):=\sum_{z\in\Lambda_{N}}b_{z}^{N,\ell,L}\bigl(\sigma_{z+1}-\sigma_{z}\bigr).$
Then Lemma 8 with $W=W_{N,\ell,L}$ yields that
$\mathbb{E}_{f_{N,0}}\left[\left(\int_{0}^{T}W_{N,\ell,L}(\sigma(t))\,dt\right)^{2}\right]\leq C\,T\,\|W_{N,\ell,L}\|_{\mathcal{H}^{-1}_{N}}^{2},$
Then by the variational characterization (25) of the $\mathcal{H}^{-1}$-norm and Young’s inequality,
$\|W_{N,\ell,L}\|_{\mathcal{H}^{-1}_{N}}^{2}\leq\frac{C}{N^{2}}\sum_{z\in\Lambda_{N}}|b_{z}^{N,\ell,L}|^{2}\leq C_{J}\,\frac{L}{N^{3}}.$
Hence
$\mathbb{E}_{f_{N,0}}\left[\left(\int_{0}^{T}W_{N,\ell,L}(\sigma(t))\,dt\right)^{2}\right]\leq C_{J}\,T\,\frac{L}{N^{3}}.$
In particular, since $L/N\to 0$,
$\lim_{\gamma\downarrow 0}\mathbb{E}_{f_{N,0}}\left|\int_{0}^{T}\frac{1}{N}\sum_{x\in\Lambda_{N}}J\!\left(\frac{x}{N}\right)\Bigl(m_{x}^{\ell}(\sigma(t))-\widetilde{m}_{x}^{\ell,L}(\sigma(t))\Bigr)\,dt\right|=0.$
Step 2: from the sliding average to the larger block average. By elementary averaging,
$\widetilde{m}_{x}^{\ell,L}=\frac{1}{2L+1}\sum_{|y|\leq L}\frac{1}{2\ell+1}\sum_{|u|\leq\ell}\sigma_{x+y+u}=\sum_{v\in\mathbb{Z}}w_{\ell,L}(v)\sigma_{x+v},$
where $w_{\ell,L}$ is a probability kernel supported in $\{-L-\ell,\dots,L+\ell\}$. On the other hand,
$m_{x}^{L}=\sum_{|v|\leq L}\frac{1}{2L+1}\sigma_{x+v}.$
A direct comparison of the weights yields
$\sup_{\sigma\in\Sigma_{N}}\sup_{x\in\Lambda_{N}}|\widetilde{m}_{x}^{\ell,L}(\sigma)-m_{x}^{L}(\sigma)|\leq C\,\frac{\ell}{L}.$
Therefore
$\left|\int_{0}^{T}\frac{1}{N}\sum_{x\in\Lambda_{N}}J\!\left(\frac{x}{N}\right)\Bigl(\widetilde{m}_{x}^{\ell,L}(\sigma(t))-m_{x}^{L}(\sigma(t))\Bigr)\,dt\right|\leq CT\|J\|_{\infty}\frac{\ell}{L}.$
Since $\ell/L\to 0$, the result follows. ∎
Corollary 7 (Comparison of local equilibrium averages at two block scales).
Assume, in addition, that the family $\Phi_{\Psi,\ell}^{\gamma}$ satisfies:
  1. (i)
    for every $\ell$, the map $m\mapsto\Phi_{\Psi,\ell}^{\gamma}(m)$ is Lipschitz on [1,1][-1,1], with Lipschitz constant bounded uniformly in $\ell$ and $\gamma$;
  2. (ii)
    there exists a constitutive law $F_{\Psi}^{\gamma}:[-1,1]\to\mathbb{R}$ such that
    $\sup_{|m|\leq 1}\bigl|\Phi_{\Psi,\ell}^{\gamma}(m)-F_{\Psi}^{\gamma}(m)\bigr|\to 0\qquad\text{as }\ell\to\infty,\ \gamma\ell\to 0.$
Then, with the same choice of $\ell_{\gamma},L_{\gamma}$ as in Theorem 6,
$\lim_{\gamma\downarrow 0}\mathbb{E}_{f_{N,0}}\left|\int_{0}^{T}\frac{1}{N}\sum_{x\in\Lambda_{N}}J\!\left(\frac{x}{N}\right)\Bigl(\Phi_{\Psi,\ell_{\gamma}}^{\gamma}(m_{x}^{\ell_{\gamma}}(\sigma(t)))-\Phi_{\Psi,L_{\gamma}}^{\gamma}(m_{x}^{L_{\gamma}}(\sigma(t)))\Bigr)\,dt\right|=0.$
Proof.
Write
$\Phi_{\Psi,\ell}^{\gamma}(m_{x}^{\ell})-\Phi_{\Psi,L}^{\gamma}(m_{x}^{L})=\Bigl[\Phi_{\Psi,\ell}^{\gamma}(m_{x}^{\ell})-\Phi_{\Psi,\ell}^{\gamma}(m_{x}^{L})\Bigr]+\Bigl[\Phi_{\Psi,\ell}^{\gamma}(m_{x}^{L})-F_{\Psi}^{\gamma}(m_{x}^{L})\Bigr]+\Bigl[F_{\Psi}^{\gamma}(m_{x}^{L})-\Phi_{\Psi,L}^{\gamma}(m_{x}^{L})\Bigr].$
The first term is controlled by the Lipschitz bound and Theorem 6; the last two terms vanish by (ii). ∎
4.3 Kac-scale regularity of mesoscopic block averages
Now we consider the observable $\Psi=d_{0}^{2}=(\eta_{0}-\eta_{1})^{2}$ in the drift term. Fix a block
$\Lambda_{L}:=\{-L,\dots,L\},\quad|\Lambda_{L}|=2L+1,$
and a block magnetization average $m=\frac{1}{2L+1}\sum_{u\in\Lambda_{L}}\eta_{u}\in[-1,1].$
Define the auxiliary canonical measure
$\nu_{L,m}^{\mathrm{aux}}(d\eta):=\frac{1}{Z_{L,m}^{\mathrm{aux}}}\mathbf{1}_{\left\{\sum_{u\in\Lambda_{L}}\eta_{u}=m(2L+1)\right\}}\prod_{u\in\Lambda_{L}}\delta_{\{-1,1\}}(d\eta_{u}).$
Now we give the local equilibrium average for the observable $\Psi=d_{0}^{2}$.
Lemma 10.
There exists a constant $C_{\beta}>0$, depending only on $\beta$ and $\|\mathfrak{K}^{\prime}\|_{L^{\infty}}$, such that
$\sup_{|m|\leq 1}\left|\Phi_{\Psi,L}^{\gamma}(m)-\frac{2(2L+1)}{2L}(1-m^{2})\right|\leq C_{\beta}\,\gamma^{2}L^{3}.$
In particular, if $\gamma^{2}L^{3}\to 0$, then $\Phi_{\Psi,L}^{\gamma}(m)=a_{0,L,\gamma}+a_{2,L,\gamma}m^{2}+\rho_{\gamma,L}(m),$ with
$a_{0,L,\gamma}=2+\frac{1}{L},\qquad a_{2,L,\gamma}=-2-\frac{1}{L},\qquad\sup_{|m|\leq 1}|\rho_{\gamma,L}(m)|\leq C_{\beta}\,\gamma^{2}L^{3}.$
Proof.
Since $d_{0}^{2}=(\eta_{0}-\eta_{1})^{2}=2(1-\eta_{0}\eta_{1}),$ it suffices to compute the two-point function.
Under $\nu_{L,m}^{\mathrm{aux}}$, exchangeability implies that for every $i\neq j$, $\mathbb{E}_{\nu_{L,m}^{\mathrm{aux}}}[\eta_{i}\eta_{j}]=:c_{L}(m).$ Now
$\left(\sum_{u\in\Lambda_{L}}\eta_{u}\right)^{2}=\sum_{u\in\Lambda_{L}}\eta_{u}^{2}+\sum_{u\neq v}\eta_{u}\eta_{v}=(2L+1)+2L(2L+1)c_{L}(m).$
Since $\sum_{u\in\Lambda_{L}}\eta_{u}=m(2L+1)$ on the canonical sector,
$(2L+1)^{2}m^{2}=(2L+1)+2L(2L+1)c_{L}(m),$
hence
$c_{L}(m)=\frac{(2L+1)m^{2}-1}{2L}.$
Therefore
$\int d_{0}^{2}\,\nu_{L,m}^{\mathrm{aux}}(d\eta)=2\bigl(1-c_{L}(m)\bigr)=\frac{2L+1}{L}(1-m^{2}).$
We now compare $\nu_{L,m}^{\mathrm{aux}}$ with the true Ising–Kac local canonical Gibbs measure. Write
$H_{\gamma,L}^{\mathrm{blk}}(\eta)=-\frac{1}{2}\sum_{u,v\in\Lambda_{L}}\kappa_{\gamma}(u-v)\eta_{u}\eta_{v}.$
Set $\bar{\kappa}_{\gamma,L}:=\gamma\,\mathfrak{K}(0).$ Then
$H_{\gamma,L}^{\mathrm{blk}}(\eta)=-\frac{1}{2}\bar{\kappa}_{\gamma,L}\left(\sum_{u\in\Lambda_{L}}\eta_{u}\right)^{2}-\frac{1}{2}\sum_{u,v\in\Lambda_{L}}\bigl(\kappa_{\gamma}(u-v)-\bar{\kappa}_{\gamma,L}\bigr)\eta_{u}\eta_{v}.$
On the fixed-magnetization sector
$\Omega_{L,m}:=\left\{\eta\in\{-1,1\}^{\Lambda_{L}}:\ \sum_{u\in\Lambda_{L}}\eta_{u}=m(2L+1)\right\},$
the first term is constant. Hence the configuration dependence is entirely contained in
$R_{\gamma,L}(\eta):=-\frac{1}{2}\sum_{u,v\in\Lambda_{L}}\bigl(\kappa_{\gamma}(u-v)-\bar{\kappa}_{\gamma,L}\bigr)\eta_{u}\eta_{v}.$
By the mean value theorem and $\mathfrak{K}\in W^{1,\infty}$,
$|\kappa_{\gamma}(u-v)-\gamma\mathfrak{K}(0)|=\gamma\bigl|\mathfrak{K}(\gamma(u-v))-\mathfrak{K}(0)\bigr|\leq\gamma^{2}\|\mathfrak{K}^{\prime}\|_{L^{\infty}}|u-v|.$
Therefore
$|R_{\gamma,L}(\eta)|\leq\frac{1}{2}\gamma^{2}\|\mathfrak{K}^{\prime}\|_{L^{\infty}}\sum_{u,v\in\Lambda_{L}}|u-v|\leq C\,\gamma^{2}L^{3}.$
Consequently, on $\Omega_{L,m}$,
$\frac{d\mu_{\gamma,L,m}^{\mathrm{can}}}{d\nu_{L,m}^{\mathrm{aux}}}(\eta)=\frac{e^{-\beta R_{\gamma,L}(\eta)}}{\nu_{L,m}^{\mathrm{aux}}(e^{-\beta R_{\gamma,L}})},$
and the uniform bound on $R_{\gamma,L}$ yields
$\left\|\mu_{\gamma,L,m}^{\mathrm{can}}-\nu_{L,m}^{\mathrm{aux}}\right\|_{\mathrm{TV}}\leq C_{\beta}\,\gamma^{2}L^{3}.$
Since $0\leq d_{0}^{2}\leq 4$, we obtain
$\left|\Phi_{\Psi,L}^{\gamma}(m)-\int d_{0}^{2}\,\nu_{L,m}^{\mathrm{aux}}(d\eta)\right|\leq 4\left\|\mu_{\gamma,L,m}^{\mathrm{can}}-\nu_{L,m}^{\mathrm{aux}}\right\|_{\mathrm{TV}}\leq C_{\beta}\,\gamma^{2}L^{3}.$
Combining this with the exact formula under $\nu_{L,m}^{\mathrm{aux}}$ proves the lemma. ∎
Lemma 11 (Kac-scale regularity of mesoscopic block averages).
Define
$\bar{h}_{\gamma,L_{\gamma}}(\sigma;x):=\sum_{z\in\Lambda_{N}}\kappa_{\gamma}(x-z)\,m_{z}^{L_{\gamma}}(\sigma).$
Then, for every $J\in C^{\infty}(\mathbb{T})$,
$\lim_{\gamma\downarrow 0}\mathbb{E}_{f_{N,0}}\left[\int_{0}^{T}\frac{1}{N}\sum_{x\in\Lambda_{N}}\left|J\!\left(\frac{x}{N}\right)\right|\sum_{z\in\Lambda_{N}}\kappa_{\gamma}(x-z)\bigl(m_{z}^{L_{\gamma}}(\sigma(t))-\bar{h}_{\gamma,L_{\gamma}}(\sigma(t);x)\bigr)^{2}\,dt\right]=0.$
Proof.
For brevity write
$m_{z}^{L}(t):=m_{z}^{L_{\gamma}}(\sigma(t)),\qquad\bar{h}_{\gamma,L}(t;x):=\bar{h}_{\gamma,L_{\gamma}}(\sigma(t);x).$
Set
$\mathcal{R}_{\gamma}:=\int_{0}^{T}\frac{1}{N}\sum_{x\in\Lambda_{N}}\left|J\!\left(\frac{x}{N}\right)\right|\sum_{z\in\Lambda_{N}}\kappa_{\gamma}(x-z)\bigl(m_{z}^{L}(t)-\bar{h}_{\gamma,L}(t;x)\bigr)^{2}\,dt.$
Since $\kappa_{\gamma}\geq 0$ and $\sum_{z}\kappa_{\gamma}(x-z)=1$, $\bar{h}_{\gamma,L}(t;x)$ is the $\kappa_{\gamma}(x-\cdot)$-weighted mean of $\{m_{z}^{L}(t)\}_{z}$. Hence, by the minimizing property of the weighted mean,
$\sum_{z\in\Lambda_{N}}\kappa_{\gamma}(x-z)\bigl(m_{z}^{L}(t)-\bar{h}_{\gamma,L}(t;x)\bigr)^{2}\leq\sum_{z\in\Lambda_{N}}\kappa_{\gamma}(x-z)\bigl(m_{z}^{L}(t)-m_{x}^{L}(t)\bigr)^{2}.$
Since $|m_{z}^{L}(t)|\leq 1$,
$\bigl(m_{z}^{L}(t)-m_{x}^{L}(t)\bigr)^{2}\leq 2\,|m_{z}^{L}(t)-m_{x}^{L}(t)|.$
Therefore
$\mathcal{R}_{\gamma}\leq 2\int_{0}^{T}\frac{1}{N}\sum_{x\in\Lambda_{N}}\left|J\!\left(\frac{x}{N}\right)\right|\sum_{y\in\Lambda_{N}}\kappa_{\gamma}(y)\,|m_{x+y}^{L}(t)-m_{x}^{L}(t)|\,dt.$
Fix $A>0$, and define $M_{\gamma}^{A}:=\left\lfloor\frac{A}{\gamma}\right\rfloor.$ Split the Kac sum as
$\sum_{y}\kappa_{\gamma}(y)|m_{x+y}^{L}-m_{x}^{L}|=\sum_{|y|\leq M_{\gamma}^{A}}\kappa_{\gamma}(y)|m_{x+y}^{L}-m_{x}^{L}|+\sum_{|y|>M_{\gamma}^{A}}\kappa_{\gamma}(y)|m_{x+y}^{L}-m_{x}^{L}|.$
and
$\mathcal{R}_{\gamma}\leq 2\mathcal{R}_{\gamma,A}^{\mathrm{near}}+2\mathcal{R}_{\gamma,A}^{\mathrm{tail}}.$
For the near-field part, since $\mathfrak{K}\in L^{\infty}$, $\sup_{y}\kappa_{\gamma}(y)\leq\gamma\|\mathfrak{K}\|_{L^{\infty}}.$ Hence
$\mathcal{R}_{\gamma,A}^{\mathrm{near}}\leq C_{A}\int_{0}^{T}\frac{1}{N}\sum_{x\in\Lambda_{N}}\frac{1}{2M_{\gamma}^{A}+1}\sum_{|y|\leq M_{\gamma}^{A}}\left|J\!\left(\frac{x}{N}\right)\right||m_{x+y}^{L_{\gamma}}(\sigma(t))-m_{x}^{L_{\gamma}}(\sigma(t))|\,dt.$
Since
$M_{\gamma}^{A}\to\infty,\qquad\frac{L_{\gamma}}{M_{\gamma}^{A}}\sim\frac{\gamma L_{\gamma}}{A}\to 0,\qquad\frac{M_{\gamma}^{A}}{N}\sim\frac{A}{\gamma N}\to 0,$
Theorem 6 applies with small block size $L_{\gamma}$ and larger averaging window $M_{\gamma}^{A}$, so for every fixed $A>0$,
$\lim_{\gamma\downarrow 0}\mathbb{E}_{f_{N,0}}[\mathcal{R}_{\gamma,A}^{\mathrm{near}}]=0.$
For the tail part, since $|m_{x+y}^{L}-m_{x}^{L}|\leq 2$,
$\mathcal{R}_{\gamma,A}^{\mathrm{tail}}\leq 2T\|J\|_{\infty}\sum_{|y|>M_{\gamma}^{A}}\kappa_{\gamma}(y).$
Using $\kappa_{\gamma}(y)=\gamma\mathfrak{K}(\gamma y)$ and $\mathfrak{K}\in L^{1}$,
$\limsup_{\gamma\downarrow 0}\sum_{|y|>M_{\gamma}^{A}}\kappa_{\gamma}(y)\leq C\int_{|u|>A/2}|\mathfrak{K}(u)|\,du,$
which tends to 0 as $A\to\infty$.
Combining the near-field and tail bounds yields $\lim_{\gamma\downarrow 0}\mathbb{E}_{f_{N,0}}[\mathcal{R}_{\gamma}]=0.$
We need the following basic estimates for $X_{\gamma}$.
Lemma 12.
Under scale assumption, the discrete random field $X_{\gamma}$ satisfies
$\left\|\nabla_{\varepsilon_{\gamma}}X_{\gamma}\right\|_{L^{\infty}}\leq C\,\frac{\gamma}{\delta_{\gamma}\varepsilon_{\gamma}},$ (27)
and
$\sup_{\gamma\downarrow 0}\mathbb{E}_{f_{N,0}}\left[\int_{0}^{T}\left\|X_{\gamma}\right\|_{L^{p}}dt\right]<\infty.$ (28)
Proof.
Since $X_{\gamma}(t,\varepsilon_{\gamma}i)=\delta_{\gamma}^{-1}\sum_{j\in\Lambda_{N}}\kappa_{\gamma}(i-j)\sigma_{j}\big(\frac{t}{\alpha_{\gamma}}\big),$ we have
$\left|\nabla_{\varepsilon_{\gamma}}X_{\gamma}(t,\varepsilon_{\gamma}i)\right|=\frac{1}{\delta_{\gamma}\varepsilon_{\gamma}}\left|h_{\gamma}\Big(\sigma\big(\frac{t}{\alpha_{\gamma}}\big);i+1\Big)-h_{\gamma}\Big(\sigma\big(\frac{t}{\alpha_{\gamma}}\big);i\Big)\right|.$
By $\kappa_{\gamma}(z)=\gamma\,\mathfrak{K}(\gamma z),\mathfrak{K}\in W^{1,\infty}(\mathbb{R}),$ and $|\sigma(j)|\leq 1$, we obtain
$\left|h_{\gamma}(\sigma;i+1)-h_{\gamma}(\sigma;i)\right|\leq\sum_{j\in\Lambda_{N}}\bigl|\kappa_{\gamma}(i+1-j)-\kappa_{\gamma}(i-j)\bigr|\leq C\,\gamma.$
Therefore
$\sup_{\gamma\downarrow 0}\mathbb{E}_{f_{N,0}}\left[\int_{0}^{T}\Big(\varepsilon_{\gamma}\sum_{i\in\Lambda_{N}}\Big|X_{\gamma}(t,\varepsilon_{\gamma}i)\Big|^{p}\Big)^{\frac{1}{p}}dt\right]\leq\sup_{\gamma\downarrow 0}\mathbb{E}_{f_{N,0}}\left[\int_{0}^{T}\sup_{i\in\Lambda_{N}}\Big|X_{\gamma}(t,\varepsilon_{\gamma}i)\Big|dt\right]<\infty.$
this becomes estimate (28) ∎
Now we derive the nonlinear drift for $\Psi=d_{0}^{2}$ from the the second-order Boltzmann–Gibbs principle.
Proposition 8.
Assume that the nonlinear drift tested against $\phi$ satisfies
$\mathcal{D}_{\gamma}^{\mathrm{nl}}(\phi;t)=-\frac{\beta}{4}\,\frac{\varepsilon_{\gamma}^{3}}{\alpha_{\gamma}}\sum_{i\in\Lambda_{N}}d_{i}^{2}(t)\nabla_{\varepsilon_{\gamma}}X_{\gamma}(t,\varepsilon_{\gamma}i)\,\nabla_{\varepsilon_{\gamma}}(\mathcal{K}_{\gamma}^{\varepsilon}\phi)(\varepsilon_{\gamma}i)+r_{\gamma}^{(0)}(\phi;t),$
Then
$\lim_{\gamma\downarrow 0}\mathbb{E}_{f_{N,0}}\left|\int_{0}^{T}\left[\mathcal{D}_{\gamma}^{\mathrm{nl}}(\phi;t)+\frac{\beta}{4}\frac{\varepsilon_{\gamma}^{3}}{\alpha_{\gamma}}\sum_{i\in\Lambda_{N}}\Bigl(a_{0,\gamma}+a_{2,\gamma}\delta_{\gamma}^{2}X_{\gamma}(t,\varepsilon_{\gamma}i)^{2}\Bigr)\nabla_{\varepsilon_{\gamma}}X_{\gamma}(t,\varepsilon_{\gamma}i)\,\nabla_{\varepsilon_{\gamma}}(\mathcal{K}_{\gamma}^{\varepsilon}\phi)(\varepsilon_{\gamma}i)\right]dt\right|=0,$
where $a_{0,\gamma}:=2+\frac{1}{L_{\gamma}}\to 2,\quad a_{2,\gamma}:=-2-\frac{1}{L_{\gamma}}\to-2$.
Proof.
By Theorem 5, Corollary 7, and Lemma 10, the local observable $d_{i}^{2}$ can be replaced by the mesoscopic local equilibrium average By Lemma 10,
$\Phi_{\Psi,L_{\gamma}}^{\gamma}\bigl(m_{i}^{L_{\gamma}}\Big(\sigma\big(\frac{t}{\alpha_{\gamma}}\big)\Big)\bigr)=a_{0,\gamma}+a_{2,\gamma}\big(m_{i}^{L_{\gamma}}\Big(\sigma\big(\frac{t}{\alpha_{\gamma}}\big)\Big)\big)2+\rho_{\gamma,L_{\gamma}}(m_{i}^{L_{\gamma}}\Big(\sigma\big(\frac{t}{\alpha_{\gamma}}\big)\Big)),$
with
$a_{0,\gamma}=2+\frac{1}{L_{\gamma}},\qquad a_{2,\gamma}=-2-\frac{1}{L_{\gamma}},\qquad\sup_{|m|\leq 1}|\rho_{\gamma,L_{\gamma}}(m)|\leq C_{\beta}\,\gamma^{2}L_{\gamma}^{3}.$
Substituting this into the definition of $\mathcal{D}_{\gamma}^{\mathrm{nl}}(\phi;t)$ gives
$\mathcal{D}_{\gamma}^{\mathrm{nl}}(\phi;t)=-\frac{\beta}{4}\frac{\varepsilon_{\gamma}^{3}}{\alpha_{\gamma}}\sum_{i\in\Lambda_{N}}\Bigl(a_{0,\gamma}+a_{2,\gamma}(m_{i}^{L_{\gamma}}\Big(\sigma\big(\frac{t}{\alpha_{\gamma}}\big)\Big))^{2}\Bigr)\nabla_{\varepsilon_{\gamma}}X_{\gamma}(t)\,\nabla_{\varepsilon_{\gamma}}(\mathcal{K}_{\gamma}^{\varepsilon}\phi)+r_{\gamma}^{(1)}(\phi;t)+r_{\gamma}^{(2)}(\phi;t),$
where
$\lim_{\gamma\downarrow 0}\mathbb{E}_{f_{N,0}}\left|\int_{0}^{T}r_{\gamma}^{(1)}(\phi;t)\,dt\right|=0,$
and
$r_{\gamma}^{(2)}(\phi;t):=-\frac{\beta}{4}\frac{\varepsilon_{\gamma}^{3}}{\alpha_{\gamma}}\sum_{i\in\Lambda_{N}}\rho_{\gamma,L_{\gamma}}\bigl(m_{i}^{L_{\gamma}}\Big(\sigma\big(\frac{t}{\alpha_{\gamma}}\big)\Big)\bigr)\,\nabla_{\varepsilon_{\gamma}}X_{\gamma}(t,\varepsilon_{\gamma}i)\,\nabla_{\varepsilon_{\gamma}}(\mathcal{K}_{\gamma}^{\varepsilon}\phi)(\varepsilon_{\gamma}i).$
Recall that
$\bar{h}_{\gamma,L_{\gamma}}(\sigma;i):=\sum_{z\in\Lambda_{N}}\kappa_{\gamma}(i-z)m_{z}^{L_{\gamma}}(\sigma),\qquad h_{\gamma}(\sigma;i)=\delta_{\gamma}X_{\gamma}(\varepsilon_{\gamma}i).$
Combining with $|m_{i}^{L_{\gamma}}|,|\bar{h}_{\gamma,L_{\gamma}}(i)|,|h_{\gamma}(i)|\leq 1$, we have
$(m_{i}^{L_{\gamma}})^{2}-h_{\gamma}(i)^{2}=$ $\bigl((m_{i}^{L_{\gamma}})^{2}-\bar{h}_{\gamma,L_{\gamma}}(i)^{2}\bigr)+\bigl(\bar{h}_{\gamma,L_{\gamma}}(i)^{2}-h_{\gamma}(i)^{2}\bigr)$
$\leq$ $2|m_{i}^{L_{\gamma}}-\bar{h}_{\gamma,L_{\gamma}}(i)|+2|\bar{h}_{\gamma,L_{\gamma}}(i)-h_{\gamma}(i)|.$
Now we decompose
$\left|\int_{0}^{T}\frac{\varepsilon_{\gamma}^{3}}{\alpha_{\gamma}}\sum_{i}\Bigl[(m_{i}^{L_{\gamma}})^{2}-h_{\gamma}(i)^{2}\Bigr]\nabla_{\varepsilon_{\gamma}}X_{\gamma}(t,\varepsilon_{\gamma}i)\,\nabla_{\varepsilon_{\gamma}}(\mathcal{K}_{\gamma}^{\varepsilon}\phi)(\varepsilon_{\gamma}i)dt\right|\leq 2\mathcal{E}_{\gamma}^{(1)}+2\mathcal{E}_{\gamma}^{(2)},$
where
$\mathcal{E}_{\gamma}^{(1)}:=\int_{0}^{T}\frac{\varepsilon_{\gamma}^{3}}{\alpha_{\gamma}}\sum_{i\in\Lambda_{N}}|m_{i}^{L_{\gamma}}(\sigma\big(\frac{t}{\alpha_{\gamma}}\big))-\bar{h}_{\gamma,L_{\gamma}}(\sigma\big(\frac{t}{\alpha_{\gamma}}\big);i)|\,|\nabla_{\varepsilon_{\gamma}}X_{\gamma}(t,\varepsilon_{\gamma}i)\,\nabla_{\varepsilon_{\gamma}}(\mathcal{K}_{\gamma}^{\varepsilon}\phi)(\varepsilon_{\gamma}i)|\,dt,$
and
$\mathcal{E}_{\gamma}^{(2)}:=\int_{0}^{T}\frac{\varepsilon_{\gamma}^{3}}{\alpha_{\gamma}}\sum_{i\in\Lambda_{N}}\Big|\bar{h}_{\gamma,L_{\gamma}}\Big(\sigma\big(\frac{t}{\alpha_{\gamma}}\big);i\Big)-h_{\gamma}\Big(\sigma\big(\frac{t}{\alpha_{\gamma}}\big);i\Big)\Big||\nabla_{\varepsilon_{\gamma}}X_{\gamma}(t,\varepsilon_{\gamma}i)\nabla_{\varepsilon_{\gamma}}(\mathcal{K}_{\gamma}^{\varepsilon}\phi)(\varepsilon_{\gamma}i)|\,dt.$
We first treat $\mathcal{E}_{\gamma}^{(1)}$. The definition of $\mathcal{K}_{\gamma}^{\varepsilon}\phi$ implies that
$\left\|\nabla_{\varepsilon_{\gamma}}(\mathcal{K}_{\gamma}^{\varepsilon}\phi)\right\|_{L^{\infty}}\leq\varepsilon_{\gamma}\left\|\nabla_{\varepsilon_{\gamma}}\phi\right\|_{L^{\infty}}\leq C_{\phi}\,\varepsilon_{\gamma}.$ (29)
By estimate 27, we have
$\sup_{i\in\Lambda_{N}}|\nabla_{\varepsilon_{\gamma}}X_{\gamma}(t,\varepsilon_{\gamma}i)\,\nabla_{\varepsilon_{\gamma}}(\mathcal{K}_{\gamma}^{\varepsilon}\phi)(\varepsilon_{\gamma}i)|\leq\left\|\nabla_{\varepsilon_{\gamma}}X_{\gamma}\right\|_{L^{\infty}}\leq C\,\frac{\gamma}{\delta_{\gamma}\varepsilon_{\gamma}}.$ (30)
Since $\kappa_{\gamma}(0)=\gamma\mathfrak{K}(0)>0$,
$\kappa_{\gamma}(0)\,|m_{i}^{L_{\gamma}}-\bar{h}_{\gamma,L_{\gamma}}(i)|^{2}\leq\sum_{z\in\Lambda_{N}}\kappa_{\gamma}(i-z)\bigl(m_{z}^{L_{\gamma}}-\bar{h}_{\gamma,L_{\gamma}}(i)\bigr)^{2}.$
Thus
$|m_{i}^{L_{\gamma}}-\bar{h}_{\gamma,L_{\gamma}}(i)|\leq\frac{1}{\sqrt{\gamma\mathfrak{K}(0)}}\left(\sum_{z\in\Lambda_{N}}\kappa_{\gamma}(i-z)\bigl(m_{z}^{L_{\gamma}}-\bar{h}_{\gamma,L_{\gamma}}(i)\bigr)^{2}\right)^{1/2}.$
Then by Jensen’s inequality,
$\mathcal{E}_{\gamma}^{(1)}\leq$ $C_{\phi}\frac{\varepsilon_{\gamma}^{3}}{\alpha_{\gamma}}\sum_{i\in\Lambda_{N}}\frac{\gamma}{\delta_{\gamma}}\frac{1}{\sqrt{\gamma\mathfrak{K}(0)}}\left(\sum_{z}\kappa_{\gamma}(i-z)\bigl(m_{z}^{L_{\gamma}}-\bar{h}_{\gamma,L_{\gamma}}(i)\bigr)^{2}\right)^{1/2}$
$\leq$ $C_{\phi}\frac{\varepsilon_{\gamma}^{2}\gamma^{1/2}}{\alpha_{\gamma}\delta_{\gamma}}\,\frac{1}{N}\sum_{i\in\Lambda_{N}}\left(\sum_{z}\kappa_{\gamma}(i-z)\bigl(m_{z}^{L_{\gamma}}-\bar{h}_{\gamma,L_{\gamma}}(i)\bigr)^{2}\right)^{1/2}$
$\leq$ $C_{\phi}\frac{\varepsilon_{\gamma}^{2}\gamma^{1/2}}{\alpha_{\gamma}\delta_{\gamma}}\left[\frac{1}{N}\sum_{i}\sum_{z}\kappa_{\gamma}(i-z)\bigl(m_{z}^{L_{\gamma}}-\bar{h}_{\gamma,L_{\gamma}}(i)\bigr)^{2}\right]^{1/2}.$ (31)
Hence, by Lemma 11 with $J\equiv 1$, we have $\lim_{\gamma\downarrow 0}\mathbb{E}_{f_{N,0}}[\mathcal{E}_{\gamma}^{(1)}]=0.$
We next treat $\mathcal{E}_{\gamma}^{(2)}$. By the deterministic kernel matching estimate,
$\sup_{i\in\Lambda_{N}}|\bar{h}_{\gamma,L_{\gamma}}(\sigma;i)-h_{\gamma}(\sigma;i)|\leq\|\kappa_{\gamma}*\chi_{L_{\gamma}}-\kappa_{\gamma}\|_{\ell^{1}(\Lambda_{N})}\leq C\,\gamma L_{\gamma},$
where $\chi_{L_{\gamma}}(r):=\frac{1}{2L_{\gamma}+1}\mathbf{1}_{\{|r|\leq L_{\gamma}\}}.$ Using the bound (30) again, we have
$\mathcal{E}_{\gamma}^{(2)}\leq C_{\phi}\frac{\varepsilon_{\gamma}^{3}}{\alpha_{\gamma}}N\frac{\gamma}{\delta_{\gamma}}(\gamma L_{\gamma})=C_{\phi}\frac{\varepsilon_{\gamma}^{2}\gamma^{2}L_{\gamma}}{\alpha_{\gamma}\delta_{\gamma}}.$
Since $\frac{\varepsilon_{\gamma}^{2}\gamma^{2}L_{\gamma}}{\alpha_{\gamma}\delta_{\gamma}}=\frac{\varepsilon_{\gamma}^{2}\gamma^{1/2}}{\alpha_{\gamma}\delta_{\gamma}}\,\gamma^{3/2}L_{\gamma}$ and $\gamma L_{\gamma}\to 0$, we have $\gamma^{3/2}L_{\gamma}\to 0$, we have $\lim_{\gamma\downarrow 0}\mathcal{E}_{\gamma}^{(2)}=0.$ Combining with the convergences of $\mathcal{E}_{\gamma}^{(1)}$ and $\mathcal{E}_{\gamma}^{(2)}$, we obtain
$\lim_{\gamma\downarrow 0}\mathbb{E}_{f_{N,0}}\left|\int_{0}^{T}\frac{\varepsilon_{\gamma}^{3}}{\alpha_{\gamma}}\sum_{i\in\Lambda_{N}}\Bigl[(m_{i}^{L_{\gamma}}(\sigma(t)))^{2}-\delta_{\gamma}^{2}X_{\gamma}(t,\varepsilon_{\gamma}i)^{2}\Bigr]\nabla_{\varepsilon_{\gamma}}X_{\gamma}(t,\varepsilon_{\gamma}i)\,\nabla_{\varepsilon_{\gamma}}(\mathcal{K}_{\gamma}^{\varepsilon}\phi)(\varepsilon_{\gamma}i)dt\right|=0.$ (32)
It remains to control $r_{\gamma}^{(2)}$. Since $|\Lambda_{N}|=N=\varepsilon_{\gamma}^{-1}$, estimates (30) yields that
$\frac{\varepsilon_{\gamma}^{3}}{\alpha_{\gamma}}\sum_{i\in\Lambda_{N}}|\nabla_{\varepsilon_{\gamma}}X_{\gamma}(t,\varepsilon_{\gamma}i)\,\nabla_{\varepsilon_{\gamma}}(\mathcal{K}_{\gamma}^{\varepsilon}\phi)(\varepsilon_{\gamma}i)|\leq C_{\phi}\,\frac{\varepsilon_{\gamma}^{3}}{\alpha_{\gamma}}\,N\,\frac{\gamma}{\delta_{\gamma}}=C_{\phi}\,\frac{\gamma\,\varepsilon_{\gamma}^{2}}{\alpha_{\gamma}\delta_{\gamma}}.$
Using the uniform bound $\sup_{\gamma\downarrow 0}\frac{\gamma\,\varepsilon_{\gamma}^{2}}{\alpha_{\gamma}\delta_{\gamma}}<\infty$ and $\sup_{|m|\leq 1}|\rho_{\gamma,L_{\gamma}}(m)|\leq C_{\beta}\gamma^{2}L_{\gamma}^{3}$, we obtain
$\mathbb{E}_{f_{N,0}}\left[\int_{0}^{T}|r_{\gamma}^{(2)}(\phi;t)|\,dt\right]\leq C_{\phi,\beta,T}\,\gamma^{2}L_{\gamma}^{3}\sup_{\gamma\downarrow 0}\frac{\gamma\,\varepsilon_{\gamma}^{2}}{\alpha_{\gamma}\delta_{\gamma}}\to 0,\quad\text{as}\quad\gamma^{2}L_{\gamma}^{3}\to 0.$
Thus
$\lim_{\gamma\downarrow 0}\mathbb{E}_{f_{N,0}}\left|\int_{0}^{T}\left[\mathcal{D}_{\gamma}^{\mathrm{nl}}(\phi;t)+\frac{\beta}{4}\frac{\varepsilon_{\gamma}^{3}}{\alpha_{\gamma}}\sum_{i\in\Lambda_{N}}\Bigl(a_{0,\gamma}+a_{2,\gamma}\delta_{\gamma}^{2}X_{\gamma}(t,\varepsilon_{\gamma}i)^{2}\Bigr)G_{\gamma,\phi}(t,i)\right]dt\right|=0.$
This proves the proposition. ∎
We now collect the previous drift and martingale identifications into a single discrete stochastic evolution for $X_{\gamma}$.
Corollary 9.
Let
$\lambda_{\gamma}:=\frac{\varepsilon_{\gamma}^{2}}{\alpha_{\gamma}},\quad A^{\prime}_{\gamma}:=\Bigl(\frac{1}{2}-\beta\kappa_{\gamma}(1)\Bigr)\lambda_{\gamma},\quad\nu_{\gamma}:=\Bigl(\frac{1}{2}-\beta\kappa_{\gamma}(1)\Bigr)\frac{\mathfrak{m}_{2,\gamma}}{2}\frac{\varepsilon_{\gamma}^{4}}{\alpha_{\gamma}\gamma^{2}},\quad\text{and}\quad A^{\prime\prime}_{\gamma}:=\frac{a_{0,\gamma}\beta}{4}\lambda_{\gamma}.$
For every smooth test function $\phi\in C^{\infty}(\mathbb{T})$, we have
$\langle X_{\gamma}(t),\phi\rangle_{\gamma}=$ $\langle X_{\gamma}(0),\phi\rangle_{\gamma}-A_{\gamma}\int_{0}^{t}\langle X_{\gamma}(s),\Delta_{\varepsilon_{\gamma}}\phi\rangle_{\gamma}\,ds-\nu_{\gamma}\int_{0}^{t}\langle X_{\gamma}(s),\Delta_{\varepsilon_{\gamma}}^{2}\phi\rangle_{\gamma}\,ds$
$\qquad+\frac{\beta\lambda_{\gamma}}{6}\int_{0}^{t}\langle X_{\gamma}(s)^{3},\Delta_{\varepsilon_{\gamma}}\phi\rangle_{\gamma}\,ds+M_{\gamma}(t,\phi)+R_{\gamma}(t,\phi),,$
where $\lim_{\gamma\downarrow 0}\mathbb{E}_{f_{N,0}}\Bigl[\sup_{t\leq T}|R_{\gamma}(t,\phi)|\Bigr]=0.$, $\nu_{\gamma\to\nu}$ $A_{\gamma}:=A^{\prime}_{\gamma}+A^{\prime\prime}_{\gamma}\to A=\lambda\frac{(1+\beta)}{2}$, and $\lambda_{\gamma}\to\lambda$. Equivalently,
$dX_{\gamma}=\Bigl[-\nu_{\gamma}\Delta_{\varepsilon_{\gamma}}^{2}X_{\gamma}-A_{\gamma}\Delta_{\varepsilon_{\gamma}}X_{\gamma}+\frac{\beta\lambda_{\gamma}}{6}\Delta_{\varepsilon_{\gamma}}(X_{\gamma}^{3})+R_{\gamma}\Bigr]dt+dM_{\gamma}.$
5 Convergence of the noise term
In this section we analyze the limit of noise term in the one-dimensional Ising–Kac–Kawasaki dynamics. Then we show that the associated discrete stochastic convolution term $Z_{\gamma}$ converge to the stochastic convolution term $Z(t):=\sqrt{\sigma_{*}}\int_{0}^{t}S(t-r)\nabla\cdot\xi(r)dr.$
5.1 Convergence of the discrete martingale
Recall that the associated Dynkin martingale is
$M_{t}^{\gamma}(\phi):=\langle X_{\gamma}(t),\phi\rangle_{\gamma}-\langle X_{\gamma}(0),\phi\rangle_{\gamma}-\int_{0}^{t}\Bigl(\alpha_{\gamma}^{-1}\mathscr{L}_{\gamma}^{N}\langle X_{\gamma},\phi\rangle_{\gamma}\Bigr)(\sigma(s/\alpha_{\gamma}))\,ds.$
where $\phi$ is the smooth test function in $C^{\infty}(\mathbb{T})$. The structure is parallel to the multiscale analysis of the drift part, but simpler: only the zeroth-order local equilibrium coefficient survives in the quadratic variation.
Theorem 10.
For every $\phi,\psi\in C^{\infty}(\mathbb{T})$, we have
$\langle M^{\gamma}(\phi),M^{\gamma}(\psi)\rangle_{t}\longrightarrow\sigma_{*}^{2}\,t\int_{\mathbb{T}}\partial_{x}\phi(x)\,\partial_{x}\psi(x)\,dx$
in $L^{1}(\mathbb{P}_{f_{N,0}})$, uniformly for $t\in[0,T]$. Equivalently, the martingale $M_{\gamma}$ converges to the divergence-type noise $\sigma_{*}\partial_{x}\xi$.
Proof.
Using the induced increment of the Kac-smoothed field (12), the jump of the field pairing is
$\delta_{i,i+1}^{\gamma}\langle X_{\gamma}(\sigma),\phi\rangle_{\gamma}$ $:=\langle X_{\gamma}(\sigma^{i,i+1}),\phi\rangle_{\gamma}-\langle X_{\gamma}(\sigma),\phi\rangle_{\gamma}$
$=\delta_{\gamma}^{-1}\varepsilon_{\gamma}\sum_{k\in\Lambda_{N}}d_{x}\Bigl(\kappa_{\gamma}(k-i-1)-\kappa_{\gamma}(k-i)\Bigr)\phi(\varepsilon_{\gamma}k)$
$=\delta_{\gamma}^{-1}\varepsilon_{\gamma}\,d_{i}\Bigl[(\mathcal{K}_{\gamma}^{\varepsilon}\phi)(\varepsilon_{\gamma}(i+1))-(\mathcal{K}_{\gamma}^{\varepsilon}\phi)(\varepsilon_{\gamma}i)\Bigr]$
$=\delta_{\gamma}^{-1}\varepsilon_{\gamma}^{2}\,d_{i}\,\nabla_{\varepsilon_{\gamma}}(\mathcal{K}_{\gamma}^{\varepsilon}\phi)(\varepsilon_{\gamma}i).$
The pure-jump martingale $M^{\gamma}(\phi)$ has quadratic variation is
$\langle M^{\gamma}(\phi)\rangle_{t}=$ $\int_{0}^{t}\alpha_{\gamma}^{-1}\sum_{i\in\Lambda_{N}}c_{i}\Big(\sigma\big(\frac{s}{\alpha_{\gamma}}\big)\Big)\bigl(\delta_{i,,i+1}^{\gamma}\langle X_{\gamma}(\sigma),\phi\rangle_{\gamma}\bigr)^{2}\,ds$
$=$ $\frac{\varepsilon_{\gamma}^{4}}{\alpha_{\gamma}\delta_{\gamma}^{2}}\int_{0}^{t}\sum_{i\in\Lambda_{N}}d^{2}_{i}\big(\frac{s}{\alpha_{\gamma}}\big)c_{i}\Big(\sigma\big(\frac{s}{\alpha_{\gamma}}\big)\Big)\Bigl[\nabla_{\varepsilon_{\gamma}}(\mathcal{K}_{\gamma}^{\varepsilon}\phi)(\varepsilon_{\gamma}i)\Bigr]^{2}\,ds,$
Similar with the second-order Boltzmann–Gibbs principle for $d^{2}_{i}$, we now consider $c_{i}(\sigma)d^{2}_{i}$.
A direct computation shows that
$\Delta_{i}H_{\gamma}(\sigma)=d_{i}(\sigma)\sum_{j\in\Lambda_{N}}\bigl(\kappa_{\gamma}(i-j)-\kappa_{\gamma}(i+1-j)\bigr)\sigma_{j}-\bigl(\kappa_{\gamma}(0)-\kappa_{\gamma}(1)\bigr)d_{i}(\sigma)^{2}.$
Hence, since $|d_{i}(\sigma)|\leq 2$ and $|\sigma_{j}|\leq 1$,
$|\Delta_{i}H_{\gamma}(\sigma)|\leq 2\sum_{j\in\Lambda_{N}}\bigl|\kappa_{\gamma}(i-j)-\kappa_{\gamma}(i+1-j)\bigr|+4\,|\kappa_{\gamma}(0)-\kappa_{\gamma}(1)|.$
Since $\kappa_{\gamma}(z)=\gamma\,\mathfrak{K}(\gamma z),\mathfrak{K}\in W^{1,1}(\mathbb{R})\cap W^{1,\infty}(\mathbb{R}),$ the discrete $L^{1}$-difference estimate and the mean-value theorem imply
$\sup_{i,\sigma}|\Delta_{i}H_{\gamma}(\sigma)|\leq C\gamma.$
Since the function $F(z):=\frac{1}{1+e^{z}}$ satisfies
$F(0)=\frac{1}{2},\quad|F^{\prime}(z)|=\frac{e^{z}}{(1+e^{z})^{2}}\leq\frac{1}{4}\quad\text{for all }z\in\mathbb{R},$
the logistic map is globally Lipschitz. Therefore
$c_{i}(\sigma)=F(\Delta_{i}H_{\gamma}(\sigma))=\frac{1}{2}+r_{i}^{\gamma}(\sigma),\qquad\sup_{i,\sigma}|r_{i}^{\gamma}(\sigma)|\leq C\gamma.$
Hence
$d_{i}(\sigma)^{2}c_{i}(\sigma)=\frac{1}{2}\,d_{i}(\sigma)^{2}+\widetilde{r}_{i}^{\gamma}(\sigma),\qquad\sup_{i,\sigma}|\widetilde{r}_{i}^{\gamma}(\sigma)|\leq C\gamma.$
Note that $\widetilde{r}_{i}^{\gamma}$ is not a strictly local observable; it will be treated as a small uniform error term.
Using the previous decomposition,
$\langle M^{\gamma}(\phi)\rangle_{t}=$ $\frac{1}{2}\,\frac{\varepsilon_{\gamma}^{4}}{\alpha_{\gamma}\delta_{\gamma}^{2}}\int_{0}^{t}\sum_{i\in\Lambda_{N}}d_{i}^{2}\!\left(\frac{s}{\alpha_{\gamma}}\right)\,\Big|\nabla_{\varepsilon_{\gamma}}(\mathcal{K}_{\gamma}^{\varepsilon}\phi)(\varepsilon_{\gamma}i)\Big|^{2}\,ds$
$\quad+\frac{\varepsilon_{\gamma}^{4}}{\alpha_{\gamma}\delta_{\gamma}^{2}}\int_{0}^{t}\sum_{i\in\Lambda_{N}}\widetilde{r}_{i}^{\gamma}\!\left(\sigma\!\left(\frac{s}{\alpha_{\gamma}}\right)\right)\,\Big|\nabla_{\varepsilon_{\gamma}}(\mathcal{K}_{\gamma}^{\varepsilon}\phi)(\varepsilon_{\gamma}i)\Big|^{2}\,ds.$
Then by estimate (29), we have
$\Big|\frac{\varepsilon_{\gamma}^{4}}{\alpha_{\gamma}\delta_{\gamma}^{2}}\int_{0}^{t}\sum_{i\in\Lambda_{N}}\widetilde{r}_{i}^{\gamma}\!\left(\sigma\!\left(\frac{s}{\alpha_{\gamma}}\right)\right)\,\Big|\nabla_{\varepsilon_{\gamma}}(\mathcal{K}_{\gamma}^{\varepsilon}\phi)(\varepsilon_{\gamma}i)\Big|^{2}\,ds\Big|\leq C_{\phi}\,t\,\gamma\,\frac{\varepsilon_{\gamma}^{3}}{\alpha_{\gamma}\delta_{\gamma}^{2}}\quad\text{as }\gamma\,\frac{\varepsilon_{\gamma}^{3}}{\alpha_{\gamma}\delta_{\gamma}^{2}}\to 0.$
By Theorem 5 and Lemma 10, the local observable $d_{i}^{2}$ may be replaced by the local equilibrium average
$\Phi_{d_{0}^{2},L_{\gamma}}^{\gamma}(m)=a_{0,\gamma}+a_{2,\gamma}m^{2}+\rho_{\gamma,L_{\gamma}}(m).$
Therefore
$\langle M^{\gamma}(\phi)\rangle_{t}=\mathcal{Q}_{\gamma}^{(0)}(\phi;t)+\mathcal{Q}_{\gamma}^{(2)}(\phi;t)+\mathcal{R}_{\gamma}^{(\rho)}(\phi;t)+o_{\gamma}(1),$
where
$\mathcal{Q}_{\gamma}^{(0)}(\phi;t):=\widetilde{a}_{0,\gamma}\,\frac{\varepsilon_{\gamma}^{4}}{\alpha_{\gamma}\delta_{\gamma}^{2}}\int_{0}^{t}\sum_{i\in\Lambda_{N}}\Big|\nabla_{\varepsilon_{\gamma}}(\mathcal{K}_{\gamma}^{\varepsilon}\phi)(\varepsilon_{\gamma}i)\Big|^{2}\,ds,$
$\mathcal{Q}_{\gamma}^{(2)}(\phi;t):=\widetilde{a}_{2,\gamma}\,\frac{\varepsilon_{\gamma}^{4}}{\alpha_{\gamma}\delta_{\gamma}^{2}}\int_{0}^{t}\sum_{i\in\Lambda_{N}}\bigl(m_{i}^{L_{\gamma}}(\sigma(s/\alpha_{\gamma}))\bigr)^{2}\Big|\nabla_{\varepsilon_{\gamma}}(\mathcal{K}_{\gamma}^{\varepsilon}\phi)(\varepsilon_{\gamma}i)\Big|^{2}\,ds,$
and
$\mathcal{R}_{\gamma}^{(\rho)}(\phi;t):=\frac{1}{2}\,\frac{\varepsilon_{\gamma}^{4}}{\alpha_{\gamma}\delta_{\gamma}^{2}}\int_{0}^{t}\sum_{i\in\Lambda_{N}}\rho_{\gamma,L_{\gamma}}\bigl(m_{i}^{L_{\gamma}}(\sigma(s/\alpha_{\gamma}))\bigr)\Big|\nabla_{\varepsilon_{\gamma}}(\mathcal{K}_{\gamma}^{\varepsilon}\phi)(\varepsilon_{\gamma}i)\Big|^{2}\,ds.$
Since $\Big|\nabla_{\varepsilon_{\gamma}}(\mathcal{K}_{\gamma}^{\varepsilon}\phi)\Big|$ is uniformly bounded, $|\mathcal{R}_{\gamma}^{(\rho)}(\phi;t)|\leq C_{\phi}\,t\,\gamma^{2}L_{\gamma}^{3}\,\frac{\varepsilon_{\gamma}^{3}}{\alpha_{\gamma}\delta_{\gamma}^{2}}.$
Because $\widetilde{\sigma}_{\gamma}^{2}\to\sigma_{*}^{2}$, the factor $\varepsilon_{\gamma}^{3}/(\alpha_{\gamma}\delta_{\gamma}^{2})$ is bounded, and since $\gamma^{2}L_{\gamma}^{3}\to 0$,
$\sup_{t\leq T}|\mathcal{R}_{\gamma}^{(\rho)}(\phi;t)|\longrightarrow 0.$
Similar with estimate (32), we have the following replacement from Kac-scale regularity:
$\sup_{t\in[0,T]}\frac{\varepsilon_{\gamma}^{4}}{\alpha_{\gamma}\delta_{\gamma}^{2}}\int_{0}^{t}\sum_{i\in\Lambda_{N}}\Bigl[(m_{i}^{L_{\gamma}})^{2}-h_{\gamma}(i)^{2}\Bigr]\Big|\nabla_{\varepsilon_{\gamma}}(\mathcal{K}_{\gamma}^{\varepsilon}\phi)(\varepsilon_{\gamma}i)\Big|^{2}ds\rightarrow 0\quad\text{in}\quad L^{1}(\mathbb{P}_{f_{N,0}}).$
Since $h_{\gamma}(\sigma;i)=\delta_{\gamma}X_{\gamma}(\varepsilon_{\gamma}i)$, this implies
$\mathcal{Q}_{\gamma}^{(2)}(\phi;t)=\widetilde{a}_{2,\gamma}\,\frac{\varepsilon_{\gamma}^{4}}{\alpha_{\gamma}}\int_{0}^{t}\sum_{i\in\Lambda_{N}}X_{\gamma}(s,\varepsilon_{\gamma}i)^{2}\Big|\nabla_{\varepsilon_{\gamma}}(\mathcal{K}_{\gamma}^{\varepsilon}\phi)(\varepsilon_{\gamma}i)\Big|^{2}ds+o_{\gamma}(1).$
By estimate (29), we have
$|\mathcal{Q}_{\gamma}^{(2)}(\phi;t)|\leq C_{\phi}\,\frac{\varepsilon_{\gamma}^{3}}{\alpha_{\gamma}}\int_{0}^{t}\varepsilon_{\gamma}\sum_{i\in\Lambda_{N}}X_{\gamma}(s,\varepsilon_{\gamma}i)^{2}\,ds+o_{\gamma}(1).$
Then by $L^{2}$ bound (28) for $X_{\gamma}$ and $\frac{\varepsilon_{\gamma}^{3}}{\alpha_{\gamma}}\to 0,$ we conclude that
$\sup_{t\leq T}\mathbb{E}_{f_{N,0}}|\mathcal{Q}_{\gamma}^{(2)}(\phi;t)|\longrightarrow 0.$
Since the integrand is deterministic, we identify
$\mathcal{Q}_{\gamma}^{(0)}(\phi;t)=t\,\widetilde{a}_{0,\gamma}\,\frac{\varepsilon_{\gamma}^{4}}{\alpha_{\gamma}\delta_{\gamma}^{2}}\sum_{i\in\Lambda_{N}}\Big|\nabla_{\varepsilon_{\gamma}}(\mathcal{K}_{\gamma}^{\varepsilon}\phi)(\varepsilon_{\gamma}i)\Big|^{2}.$
Since $\varepsilon_{\gamma}/\gamma\to 0$, the discrete Kac convolution approximates the identity at the level of gradients:
$\varepsilon_{\gamma}\sum_{i\in\Lambda_{N}}\Big|\nabla_{\varepsilon_{\gamma}}(\mathcal{K}_{\gamma}^{\varepsilon}\phi)(\varepsilon_{\gamma}i)\Big|^{2}\rightarrow\int_{\mathbb{T}}|\partial_{x}\phi(x)|^{2}\,dx.$
Therefore
$\mathcal{Q}_{\gamma}^{(0)}(\phi;t)=t\,\left(\widetilde{a}_{0,\gamma}\frac{\varepsilon_{\gamma}^{3}}{\alpha_{\gamma}\delta_{\gamma}^{2}}\right)\left(\varepsilon_{\gamma}\sum_{i\in\Lambda_{N}}|\nabla_{\varepsilon_{\gamma}}(\mathcal{K}_{\gamma}^{\varepsilon}\phi)(\varepsilon_{\gamma}i)|^{2}\right)\longrightarrow\sigma_{*}^{2}\,t\int_{\mathbb{T}}|\partial_{x}\phi(x)|^{2}\,dx.$
Combining above estimates, we obtain
$\langle M^{\gamma}(\phi)\rangle_{t}\longrightarrow\sigma_{*}^{2}\,t\int_{\mathbb{T}}|\partial_{x}\phi(x)|^{2}\,dx\quad\text{in}\quad L^{1}(\mathbb{P}_{f_{N,0}})$
uniformly for $t\in[0,T]$.
Finally, at each Kawasaki jump across the bond $(i,i+1)$, we have
$\bigl|\Delta M^{\gamma}(\phi)\bigr|=\left|\delta_{i,i+1}^{\gamma}\langle X_{\gamma},\phi\rangle_{\gamma}\right|\leq C\,\frac{\varepsilon_{\gamma}^{2}}{\delta_{\gamma}}\,\|B_{\gamma,\phi}\|_{\ell^{\infty}}\leq C_{\phi}\,\frac{\varepsilon_{\gamma}^{2}}{\delta_{\gamma}}.$ (33)
Since $\varepsilon_{\gamma}^{2}/\delta_{\gamma}\to 0$, the jumps vanish. Martingale central limit theorem yields
$M^{\gamma}(\phi)\Longrightarrow M(\phi)$
in $D([0,T],\mathbb{R})$, where $M(\phi)$ is a continuous centered Gaussian martingale with quadratic variation
$\langle M(\phi)\rangle_{t}=\sigma_{*}^{2}\,t\int_{\mathbb{T}}|\partial_{x}\phi(x)|^{2}\,dx.$
The cross-variation follows by polarization:
$\langle M^{\gamma}(\phi),M^{\gamma}(\psi)\rangle_{t}=\frac{1}{4}\Bigl(\langle M^{\gamma}(\phi+\psi)\rangle_{t}-\langle M^{\gamma}(\phi-\psi)\rangle_{t}\Bigr)\longrightarrow\sigma_{*}^{2}\,t\int_{\mathbb{T}}\partial_{x}\phi\,\partial_{x}\psi\,dx.$
This completes the proof. ∎
5.2 Convergence of the stochastic convolution in $\mathcal{C}^{\alpha}$
Recall that $Z_{\gamma}(t):=\int_{0}^{t}S_{\gamma}(t-r)\,dM_{\gamma}(r),$ where $M_{\gamma}$ is the lattice martingale field constructed above. We set $\widetilde{Z}_{\gamma}(t):=\mathrm{Ext}_{\gamma}Z_{\gamma}(t).$ For $k\in\mathcal{B}_{\gamma}$, define the Fourier mode of martingale
$\widehat{M}_{\gamma}(t,k):=M_{t}^{\gamma}(e_{-k}).$
Now we show the convergence of the Fourier-mode bracket.
Lemma 13.
Under the Assumption $({\bf A})$ , for $k,\ell\in\mathcal{B}_{\gamma}$,
$\left\langle\widehat{M}_{\gamma}(k),\overline{\widehat{M}_{\gamma}(\ell)}\right\rangle_{t}=\delta_{k,\ell}\,q_{\gamma}(k)\,t+o_{\gamma}(1)$
in $L^{1}(\mathbb{P}_{f_{N,0}})$, uniformly for $t\in[0,T]$, where $|q_{\gamma}(k)|\leq C|k|^{2}$ uniformly in $k\in\mathcal{B}_{\gamma}$, and
$q_{\gamma}(k)\longrightarrow\sigma_{*}^{2}(2\pi k)^{2}$
locally uniformly in $k\in\mathbb{Z}$.
Proof.
Applying Theorem 10 with $\phi=e_{-k},\psi=e_{\ell},$ we obtain
$\left\langle\widehat{M}_{\gamma}(k),\overline{\widehat{M}_{\gamma}(\ell)}\right\rangle_{t}=\widetilde{\sigma}_{\gamma}^{2}\,t\,\varepsilon_{\gamma}\sum_{i\in\Lambda_{N}}\nabla_{\varepsilon_{\gamma}}(\mathcal{K}_{\gamma}^{\varepsilon}e_{-k})(\varepsilon_{\gamma}i)\,\overline{\nabla_{\varepsilon_{\gamma}}(\mathcal{K}_{\gamma}^{\varepsilon}e_{-\ell})(\varepsilon_{\gamma}i)}+o_{\gamma}(1),$
where the error converges to zero in $L^{1}(\mathbb{P}_{f_{N,0}})$, uniformly in $t\in[0,T]$.
After changing of variables m=ijm=i-j, we have
$(\mathcal{K}_{\gamma}^{\varepsilon}e_{k})(\varepsilon_{\gamma}i)=\varepsilon_{\gamma}\sum_{j\in\Lambda_{N}}\kappa_{\gamma}(i-j)e^{2\pi ik\varepsilon_{\gamma}j}=\varepsilon_{\gamma}\left(\sum_{m\in\Lambda_{N}}\kappa_{\gamma}(m)e^{-2\pi ik\varepsilon_{\gamma}m}\right)e^{2\pi ik\varepsilon_{\gamma}i}.$
Applying the discrete gradient,
$\nabla_{\varepsilon_{\gamma}}(\mathcal{K}_{\gamma}^{\varepsilon}e_{k})(\varepsilon_{\gamma}i)=\varepsilon_{\gamma}\left(\sum_{m\in\Lambda_{N}}\kappa_{\gamma}(m)e^{-2\pi ik\varepsilon_{\gamma}m}\right)\frac{e_{k}(\varepsilon_{\gamma}(i+1))-e_{k}(\varepsilon_{\gamma}i)}{\varepsilon_{\gamma}}.$
Denote $\vartheta_{\gamma}(k)=\sum_{m\in\Lambda_{N}}\gamma\,\mathfrak{K}(\gamma m)e^{-2\pi ik\varepsilon_{\gamma}m}$ and $D_{\gamma}(k):=(e^{2\pi ik\varepsilon_{\gamma}}-1)/\varepsilon_{\gamma}$.
Now we show that $\vartheta_{\gamma}(k)\to 1$ locally uniformly. Set $\xi_{\gamma,k}:=2\pi k\frac{\varepsilon_{\gamma}}{\gamma}$ so that
$\vartheta_{\gamma}(k)=\sum_{m\in\Lambda_{N}}\gamma\,\mathfrak{K}(\gamma m)e^{-i\xi_{\gamma,k}(\gamma m)}.$
Since $\mathfrak{K}\in L^{1}(\mathbb{R})$ and $\int_{\mathbb{R}}\mathfrak{K}(u)\,du=1$, its Fourier transform is continuous, and
$\widehat{\mathfrak{K}}(k):=\int_{\mathbb{R}}\mathfrak{K}(u)e^{-i\xi_{\gamma,k}u}\,du\longrightarrow 1$
locally uniformly for $|k|\leq K$. Applying the standard Riemann-sum estimate for the $W^{1,1}$-function $u\mapsto\mathfrak{K}(u)e^{-i\xi_{\gamma,k}u},$ we obtain
$\left|\vartheta_{\gamma}(k)-\int_{\mathbb{R}}\mathfrak{K}(u)e^{-i\xi_{\gamma,k}u}\,du\right|\leq\gamma(\|\mathfrak{K}^{\prime}\|_{L^{1}}+|\xi_{\gamma,k}|\|\mathfrak{K}\|_{L^{1}})\leq C\gamma\bigl(1+2\pi k\frac{\varepsilon_{\gamma}}{\gamma}\bigr).$
Therefore
$\sup_{|k|\leq K}\left|\vartheta_{\gamma}(k)-1\right|\rightarrow 0\quad\text{as}\quad\frac{\varepsilon_{\gamma}}{\gamma}\to 0.$
Then
$\sup_{i,\gamma}|\nabla_{\varepsilon_{\gamma}}(\mathcal{K}_{\gamma}^{\varepsilon}e_{k})(\varepsilon_{\gamma}i)|=|D_{\gamma}(k)||\vartheta_{\gamma}(k)|\leq C_{k},$ (34)
and
$\left\langle\widehat{M}_{\gamma}(k),\overline{\widehat{M}_{\gamma}(\ell)}\right\rangle_{t}=\widetilde{\sigma}_{\gamma}^{2}\,t\,\vartheta_{\gamma}(-k)D_{\gamma}(-k)\,\overline{\vartheta_{\gamma}(-\ell)D_{\gamma}(-\ell)}\,\varepsilon_{\gamma}\sum_{i\in\Lambda_{N}}e_{-k}(\varepsilon_{\gamma}i)e_{\ell}(\varepsilon_{\gamma}i)+o_{\gamma}(1).$
By the orthogonality of discrete Fourier model,
$\varepsilon_{\gamma}\sum_{i\in\Lambda_{N}}e_{-k}(\varepsilon_{\gamma}i)e_{\ell}(\varepsilon_{\gamma}i)=\delta_{k,\ell},$
and therefore
$\left\langle\widehat{M}_{\gamma}(k),\overline{\widehat{M}_{\gamma}(\ell)}\right\rangle_{t}=\delta_{k,\ell}\,\widetilde{\sigma}_{\gamma}^{2}\,|D_{\gamma}(k)|^{2}\,|\vartheta_{\gamma}(k)|^{2}\,t+o_{\gamma}(1).$
Next, note that
$|D_{\gamma}(k)|=\left|\frac{e^{2\pi ik\varepsilon_{\gamma}}-1}{\varepsilon_{\gamma}}\right|=\frac{2|\sin(\pi k\varepsilon_{\gamma})|}{\varepsilon_{\gamma}}\leq 2\pi|k|,$
and for each fixed $k$, $D_{\gamma}(k)\longrightarrow 2\pi ik,|D_{\gamma}(k)|^{2}\longrightarrow(2\pi k)^{2}.$ Therefore
$q_{\gamma}(k)=\widetilde{\sigma}_{\gamma}^{2}\,|D_{\gamma}(k)|^{2}\,|\vartheta_{\gamma}(k)|^{2}\longrightarrow\sigma_{*}^{2}(2\pi k)^{2}$
locally uniformly in $k\in\mathbb{Z}$. This completes the proof.
Lemma 14.
Under scale assumption, the family stochastic convolution term $\{\widetilde{Z}_{\gamma}\}_{\gamma}$ has uniform bound
$\sup_{\gamma>0}\mathbb{E}\Bigl[\sup_{t\leq T}\|\widetilde{Z}_{\gamma}(t)\|_{C^{\alpha}(\mathbb{T})}\Bigr]<\infty,\quad\alpha<\frac{1}{2},$ (35)
Moreover, $\{\widetilde{Z}_{\gamma}\}_{\gamma}$ is tight in $C([0,T],\mathcal{C}^{\alpha}(\mathbb{T}))$ for every $\alpha<\frac{1}{2}$.
Proof.
By Fourier expansion, the Littlewood–Paley block of $\widetilde{Z}_{\gamma}$ is given by
$\Delta_{j}\widetilde{Z}_{\gamma}(t,x)=\sum_{k\in\mathcal{A}_{j}\cap\mathcal{B}_{\gamma}}\varrho_{q}(k)\int_{0}^{t}e^{-(t-r)\lambda_{\gamma}(k)}\,d\widehat{M}_{\gamma}(r,k)\,e_{k}(x).$
Fix $p\geq 2$. Burkholder–Davis–Gundy inequality yields
$\mathbb{E}\Bigl[\sup_{t\leq T}\|\Delta_{j}\widetilde{Z}_{\gamma}(t)\|_{L^{p}}^{p}\Bigr]\lesssim_{p}\mathbb{E}\Bigl[\Bigl(\sup_{x\in\mathbb{T}}\langle\Delta_{j}\widetilde{Z}_{\gamma}(\cdot,x)\rangle_{T}\Bigr)^{p/2}\Bigr].$
By Lemma 13, and $|\mathcal{A}_{j}|\sim 2^{q}$ and $|k|\sim 2^{q}$ on $\mathcal{A}_{j}$,
$\langle\Delta_{j}\widetilde{Z}_{\gamma}(\cdot,x)\rangle_{T}\lesssim\sum_{k\in\mathcal{A}_{j}\cap\mathcal{B}_{\gamma}}\int_{0}^{T}e^{-2(T-r)\lambda_{\gamma}(k)}\,q_{\gamma}(k)\,dr\lesssim\sum_{k\in\mathcal{A}_{j}\cap\mathcal{B}_{\gamma}}|k|^{-2}\lesssim 2^{-q}.$
Therefore
$\sup_{\gamma>0}\mathbb{E}\Bigl[\sup_{t\leq T}\|\Delta_{j}\widetilde{Z}_{\gamma}(t)\|_{L^{p}}^{p}\Bigr]\lesssim_{p}2^{-qp/2}.$
Let $\beta\in(0,\frac{1}{2})$. Then
$\sup_{\gamma>0}\mathbb{E}\Bigl[\sup_{t\leq T}\|\widetilde{Z}_{\gamma}(t)\|_{B^{\beta}_{p,p}}^{p}\Bigr]=$ $\sup_{\gamma>0}\mathbb{E}\Bigl[\sup_{t\leq T}2^{\beta pq}\|\Delta_{j}\widetilde{Z}_{\gamma}(t)\|_{L^{p}}^{p}\Bigr]$
$\leq$ $\sum_{q\geq-1}2^{\beta pq}2^{-qp/2}<\infty,$ (36)
Choosing $p$ large enough. By Besov embedding $B^{\beta}_{p,p}(\mathbb{T})\hookrightarrow\mathcal{C}^{\alpha}(\mathbb{T})$, we obtain the bound (35)
Now we show the time increments and tightness of $\{\widetilde{Z}_{\gamma}\}_{\gamma}$. For $0\leq s<t\leq T$, write
$\Delta_{j}\widetilde{Z}_{\gamma}(t)-\Delta_{j}\widetilde{Z}_{\gamma}(s)=I_{\gamma,q}^{(1)}(s,t)+I_{\gamma,q}^{(2)}(s,t),$
where
$I_{\gamma,q}^{(1)}(s,t):=\sum_{k\in\mathcal{A}_{j}\cap\mathcal{B}_{\gamma}}\varrho_{q}(k)\int_{s}^{t}e^{-(t-r)\lambda_{\gamma}(k)}\,d\widehat{M}_{\gamma}(r,k)\,e_{k},$
and
$I_{\gamma,q}^{(2)}(s,t):=\sum_{k\in\mathcal{A}_{j}\cap\mathcal{B}_{\gamma}}\varrho_{q}(k)\int_{0}^{s}\bigl(e^{-(t-r)\lambda_{\gamma}(k)}-e^{-(s-r)\lambda_{\gamma}(k)}\bigr)\,d\widehat{M}_{\gamma}(r,k)\,e_{k}.$
Applying Burkholder–Davis–Gundy inequality together with Lemma 13 and semigroup group estimates 4, we obtain
$\sup_{\gamma}\mathbb{E}\Bigl[\|\Delta_{j}(\widetilde{Z}_{\gamma}(t)-\widetilde{Z}_{\gamma}(s))\|_{L^{p}}^{p}\Bigr]\lesssim_{p}|t-s|^{\vartheta p/2}\,2^{-q(1-4\vartheta)p/2}.$
Now we choose $\beta$ with $0<\beta<\frac{1}{2}-2\vartheta,$ so that
$\sum_{q\geq-1}2^{\beta pq}2^{-q(1-4\vartheta)p/2}<\infty,$
Then
$\sup_{\gamma}\mathbb{E}\Bigl[\|\widetilde{Z}_{\gamma}(t)-\widetilde{Z}_{\gamma}(s)\|_{B^{\beta}_{p,p}}^{p}\Bigr]\lesssim|t-s|^{\vartheta p/2}.$
The Kolmogorov’s criterion yields tightness of $\{\widetilde{Z}_{\gamma}\}_{\gamma}$ in $C([0,T],B^{\beta}_{p,p}(\mathbb{T})).$ Since $B^{\beta}_{p,p}(\mathbb{T})\hookrightarrow\mathcal{C}^{\alpha}(\mathbb{T})$, the family $\{\widetilde{Z}_{\gamma}\}_{\gamma}$ is tight in $C([0,T],\mathcal{C}^{\alpha}(\mathbb{T}))$ for every $\alpha<\frac{1}{2}$. ∎
Now we show the convergence of the stochastic convolution.
Theorem 11.
For every $\alpha<\frac{1}{2}$, the discrete stochastic convolution $\widetilde{Z}_{\gamma}\Longrightarrow Z$ in $C([0,T],\mathcal{C}^{\alpha}(\mathbb{T})).$
Proof.
Lemma 14 yields that, passing to a further subsequence if necessary, we have
$\widetilde{Z}_{\gamma_{n}}\Longrightarrow\widetilde{Z}\qquad\text{in }C([0,T],B^{\beta}_{p,p}(\mathbb{T})).$
To identify the limit, it is enough to show that every convergent subsequence has limit equal to $Z$. We now prove that $\widetilde{Z}=Z$ in law.
Fix $Q>0$, and let $\mathcal{K}_{Q}:=\{k\in\mathbb{Z}:\ |k|\leq C\,2^{Q}\}.$ For $n$ sufficiently large, $\mathcal{K}_{Q}\subset\mathcal{B}_{\gamma_{n}}$. Note that for each fixed $k\in\mathbb{Z}$, the estimate (34) implies that $M_{\gamma}(\cdot,k)$ satisfies the following jump condition
$\sup_{t\leq T}|\Delta\widehat{M}_{\gamma}(t,k)|\leq 2\frac{\varepsilon_{\gamma}^{2}}{\delta_{\gamma}}\sup_{i,\gamma}|\nabla_{\varepsilon_{\gamma}}(\mathcal{K}_{\gamma}^{\varepsilon}e_{k})(\varepsilon_{\gamma}i)|\leq C_{k}\,\frac{\varepsilon_{\gamma}^{2}}{\delta_{\gamma}}\to 0.$
Then by Lemma 13 and Skorokhod’s representation theorem, after replacing the processes by copies on a common probability space, we may assume that
$\sup_{t\leq T}\max_{k\in\mathcal{K}_{Q}}\bigl|\widehat{M}_{\gamma_{n}}(t,k)-\widehat{M}(t,k)\bigr|\longrightarrow 0\qquad\text{a.s.}$
For $k\in\mathcal{K}_{Q}$, define
$\widehat{\widetilde{Z}}_{\gamma_{n}}(t,k)=\int_{0}^{t}e^{-(t-r)\lambda_{\gamma_{n}}(k)}\,d\widehat{M}_{\gamma_{n}}(r,k),\qquad\widehat{Z}(t,k)=\int_{0}^{t}e^{-(t-r)(2\pi k)^{4}}\,d\widehat{M}(r,k).$
We claim that, for each fixed $k\in\mathcal{K}_{Q}$,
$\sup_{t\leq T}\bigl|\widehat{\widetilde{Z}}_{\gamma_{n}}(t,k)-\widehat{Z}(t,k)\bigr|\longrightarrow 0\qquad\text{a.s.}$
We decompose $\widehat{\widetilde{Z}}_{\gamma_{n}}(t,k)-\widehat{Z}(t,k)=A_{n}(t,k)+B_{n}(t,k),$ where
$A_{n}(t,k):=\int_{0}^{t}\Bigl(e^{-(t-r)\lambda_{\gamma_{n}}(k)}-e^{-(t-r)(2\pi k)^{4}}\Bigr)\,d\widehat{M}_{\gamma_{n}}(r,k),$
and
$B_{n}(t,k):=\int_{0}^{t}e^{-(t-r)(2\pi k)^{4}}\,d\bigl(\widehat{M}_{\gamma_{n}}-\widehat{M}\bigr)(r,k).$
We first estimate $A_{n}$. After integration by parts, we have
$\sup_{t\leq T}|A_{n}(t,k)|\leq\sup_{r\leq T}|\widehat{M}_{\gamma_{n}}(r,k)|\int_{0}^{T}\left|\lambda_{\gamma_{n}}(k)e^{-u\lambda_{\gamma_{n}}(k)}-(2\pi k)^{4}e^{-u(2\pi k)^{4}}\right|\,du.$
Since $k$ is fixed and $\lambda_{\gamma_{n}}(k)\longrightarrow(2\pi k)^{4},$ the integral tends to 0. Moreover, since $\widehat{M}_{\gamma_{n}}(\cdot,k)\to\widehat{M}(\cdot,k)$ uniformly, $\sup_{r\leq T}|\widehat{M}_{\gamma_{n}}(r,k)|$ is almost surely bounded for large $n$. Hence
$\sup_{t\leq T}|A_{n}(t,k)|\longrightarrow 0\qquad\text{a.s.}$
We next estimate $B_{n}$. Using integration by parts, we have
$B_{n}(t,k)=\bigl(\widehat{M}_{\gamma_{n}}(t,k)-\widehat{M}(t,k)\bigr)-(2\pi k)^{4}\int_{0}^{t}e^{-(t-r)(2\pi k)^{4}}\bigl(\widehat{M}_{\gamma_{n}}(r,k)-\widehat{M}(r,k)\bigr)\,dr.$
Hence by the Skorokhod representation, as $\gamma\to 0$,
$\sup_{t\leq T}|B_{n}(t,k)|\leq\left(1+(2\pi k)^{4}T\right)\sup_{r\leq T}\bigl|\widehat{M}_{\gamma_{n}}(r,k)-\widehat{M}(r,k)\bigr|\to 0,\quad\text{a.s.}\quad$
Combining with convergences of $A_{n}(t,k)$ and $B_{n}(t,k)$, we conclude that for every $p\in[1,\infty]$ and $j\leq Q$,
$\sup_{t\leq T}\|\Delta_{j}(\widetilde{Z}_{\gamma_{n}}(t)-Z(t))\|_{L^{p}(\mathbb{T})}\leq C_{j}\max_{k\in\mathcal{A}_{j}}\sup_{t\leq T}\bigl|\widehat{\widetilde{Z}}_{\gamma_{n}}(t,k)-\widehat{Z}(t,k)\bigr|\longrightarrow 0\qquad\text{a.s.}$
Summing over $j\leq Q$, we obtain
$\sup_{t\leq T}\left\|\sum_{j\leq Q}\Delta_{j}(\widetilde{Z}_{\gamma_{n}}(t)-Z(t))\right\|_{B^{\beta}_{p,p}}\longrightarrow 0\qquad\text{a.s.}$
It remains to control the high-frequency tail. By Step 1,
$\sup_{n}\mathbb{E}\Bigl[\sup_{t\leq T}\sum_{j>Q}2^{\beta pq}\|\Delta_{j}\widetilde{Z}_{\gamma_{n}}(t)\|_{L^{p}}^{p}\Bigr]\lesssim\sum_{j>Q}2^{\beta pq}2^{-qp/2}\longrightarrow 0$
as $Q\to\infty$. The same estimate holds for $Z$, since its Fourier modes satisfy $\mathbb{E}\bigl|\widehat{Z}(t,k)\bigr|^{2}\lesssim|k|^{-2},$ and the same Littlewood–Paley argument applies. Hence, for every $\varepsilon>0$, one may first choose $Q$ so large that
$\mathbb{P}\!\left(\sup_{t\leq T}\big\|\sum_{j>Q}\Delta_{j}\widetilde{Z}_{\gamma_{n}}(t)\big\|_{B^{\beta}_{p,p}}+\sup_{t\leq T}\big\|\sum_{j>Q}\Delta_{j}Z(t)\big\|_{B^{\beta}_{p,p}}>\varepsilon\right)$
is arbitrarily small, uniformly in $n$. Then, for this fixed $Q$, the low-frequency part converges almost surely to zero. Therefore
$\sup_{t\leq T}\|\widetilde{Z}_{\gamma_{n}}(t)-Z(t)\|_{B^{\beta}_{p,p}}\longrightarrow 0\qquad\text{in probability.}$
This shows that every sub-sequential limit $\widetilde{Z}$ coincides with $Z$. Hence $\widetilde{Z}_{\gamma}\rightarrow Z$ in $C([0,T],B^{\beta}_{p,p}(\mathbb{T})).$ Finally, since $B^{\beta}_{p,p}(\mathbb{T})\hookrightarrow\mathcal{C}^{\alpha}(\mathbb{T})$ for some $\beta>\alpha+\frac{1}{p}$, we conclude $\widetilde{Z}_{\gamma}\rightarrow Z$ in $C([0,T],\mathcal{C}^{\alpha}(\mathbb{T}))$ for every $\alpha<\frac{1}{2}$. ∎
6 Energy estimates for $Y_{\gamma}$
In this section, we show the uniform discrete $H^{-1}$-energy estimate for $Y_{\gamma}$.
Proposition 12.
Assume that $Y_{\gamma}$ solves
$\partial_{t}Y_{\gamma}=-\nu_{\gamma}\Delta_{\varepsilon_{\gamma}}^{2}Y_{\gamma}-A_{\gamma}\Delta_{\varepsilon_{\gamma}}(Y_{\gamma}+Z_{\gamma})+\frac{\beta\lambda_{\gamma}}{6}\Delta_{\varepsilon_{\gamma}}\bigl((Y_{\gamma}+Z_{\gamma})^{3}\bigr)+R_{\gamma},\quad Y_{\gamma}(0)=X_{\gamma}(0),$
on the lattice subspace $\Lambda_{\varepsilon_{\gamma}}$. Then there exist constants $C_{T}>0$, independent of $\gamma$, such that for every $T>0$,
$\sup_{t\in[0,T]}\|Y_{\gamma}(t)\|_{-1,\gamma}^{2}+$ $\int_{0}^{T}\|Y_{\gamma}(s)\|_{1,\gamma}^{2}\,ds+\int_{0}^{T}\|Y_{\gamma}(s)\|_{L^{4}_{\gamma}}^{4}\,ds$
$\leq C_{T}\left(\|X_{\gamma}(0)\|_{-1,\gamma}^{2}+1+\int_{0}^{T}\|Z_{\gamma}(s)\|_{L^{4}_{\gamma}}^{4}\,ds+\int_{0}^{T}\|R_{\gamma}(s)\|_{-1,\gamma}^{2}\,ds\right).$ (37)
We also have the uniform time regularity
$\sup_{\gamma>0}\mathbb{E}\Bigl[\|\partial_{t}\widetilde{Y}_{\gamma}\|_{L^{4/3}(0,T;H^{-3})}^{4/3}\Bigr]<\infty.$ (38)
Proof.
Testing the equation against $(-\Delta_{\varepsilon_{\gamma}})^{-1}Y_{\gamma}$, we obtain
$\frac{1}{2}\nu_{\gamma}\frac{d}{dt}\|Y_{\gamma}\|_{-1,\gamma}^{2}+\|Y_{\gamma}\|_{1,\gamma}^{2}=-A_{\gamma}\langle Y_{\gamma}+Z_{\gamma},Y_{\gamma}\rangle_{\gamma}-\frac{\beta\lambda_{\gamma}}{6}\langle(Y_{\gamma}+Z_{\gamma})^{3},Y_{\gamma}\rangle_{\gamma}+\langle R_{\gamma},(-\Delta_{\varepsilon_{\gamma}})^{-1}Y_{\gamma}\rangle_{\gamma}.$
For the linear term, by the discrete Poincaré inequality and Young’s inequality,
$|\langle Y_{\gamma}+Z_{\gamma},Y_{\gamma}\rangle_{\gamma}|\leq\eta\|Y_{\gamma}\|_{1,\gamma}^{2}+C_{\eta}\|Z_{\gamma}\|_{L^{2}_{\gamma}}^{2}.$
For the nonlinear term, using
$(Y_{\gamma}+Z_{\gamma})^{3}Y_{\gamma}=Y_{\gamma}^{4}+3Y_{\gamma}^{3}Z_{\gamma}+3Y_{\gamma}^{2}Z_{\gamma}^{2}+Y_{\gamma}Z_{\gamma}^{3}$
and Young’s inequality, one obtains the pointwise lower bound
$(Y_{\gamma}+Z_{\gamma})^{3}Y_{\gamma}\geq c_{0}|Y_{\gamma}|^{4}-C|Z_{\gamma}|^{4},$
hence
$\frac{\beta\lambda_{\gamma}}{6}\langle(Y_{\gamma}+Z_{\gamma})^{3},Y_{\gamma}\rangle_{\gamma}\leq-c_{1}\|Y_{\gamma}\|_{L^{4}_{\gamma}}^{4}+C\|Z_{\gamma}\|_{L^{4}_{\gamma}}^{4}.$
Finally,
$|\langle R_{\gamma},(-\Delta_{\varepsilon_{\gamma}})^{-1}Y_{\gamma}\rangle_{\gamma}|\leq\frac{1}{2}\|Y_{\gamma}\|_{-1,\gamma}^{2}+\frac{1}{2}\|R_{\gamma}\|_{-1,\gamma}^{2}.$
Collecting the three estimates and choosing $\eta>0$ sufficiently small yields
$\frac{1}{2}\frac{d}{dt}\|Y_{\gamma}\|_{-1,\gamma}^{2}+c\|Y_{\gamma}\|_{1,\gamma}^{2}+c\|Y_{\gamma}\|_{L^{4}_{\gamma}}^{4}\leq C\Bigl(1+\|Z_{\gamma}\|_{L^{4}_{\gamma}}^{4}+\|Y_{\gamma}\|_{-1,\gamma}^{2}+\|R_{\gamma}\|_{-1,\gamma}^{2}\Bigr).$
The integrated bound (12) follows from Gronwall’s inequality.
To see the time regularity (38), let $\varphi\in H^{3}(\mathbb{T})$. Then
$\langle\partial_{t}\widetilde{Y}_{\gamma},\varphi\rangle=-\nu_{\gamma}\langle\widetilde{Y}_{\gamma},\Delta^{2}\varphi\rangle-A_{\gamma}\langle\widetilde{Y}_{\gamma}+\widetilde{Z}_{\gamma},\Delta\varphi\rangle-\frac{\beta\lambda_{\gamma}}{6}\langle(\widetilde{Y}_{\gamma}+\widetilde{Z}_{\gamma})^{3},\Delta\varphi\rangle+\langle R_{\gamma},\varphi\rangle.$
For the bi-Laplacian term,
$|\langle\widetilde{Y}_{\gamma},\Delta_{\varepsilon_{\gamma}}^{2}\varphi\rangle|=|\langle\nabla_{\varepsilon_{\gamma}}\widetilde{Y}_{\gamma},\nabla_{\varepsilon_{\gamma}}\Delta_{\varepsilon_{\gamma}}\varphi\rangle|\leq C\|\widetilde{Y}_{\gamma}\|_{H^{1}}\|\varphi\|_{H^{3}}.$
For the linear lower-order term,
$|\langle\widetilde{Y}_{\gamma}+\widetilde{Z}_{\gamma},\Delta\varphi\rangle|\leq\bigl(\|\widetilde{Y}_{\gamma}\|_{L^{2}}+\|\widetilde{Z}_{\gamma}\|_{L^{2}}\bigr)\|\varphi\|_{H^{2}}\leq C\bigl(\|\widetilde{Y}_{\gamma}\|_{H^{1}}+\|\widetilde{Z}_{\gamma}\|_{L^{\infty}}\bigr)\|\varphi\|_{H^{3}}.$
For the cubic term, since in one dimension $H^{3}(\mathbb{T})\hookrightarrow W^{2,\infty}(\mathbb{T})$,
$|\langle(\widetilde{Y}_{\gamma}+\widetilde{Z}_{\gamma})^{3},\Delta\varphi\rangle|\leq\|(\widetilde{Y}_{\gamma}+\widetilde{Z}_{\gamma})^{3}\|_{L^{1}}\,\|\Delta\varphi\|_{L^{\infty}}\leq C\|\widetilde{Y}_{\gamma}+\widetilde{Z}_{\gamma}\|_{L^{3}}^{3}\,\|\varphi\|_{H^{3}}.$
By Hölder inequality and the $L^{4}$-bound on $\widetilde{Y}_{\gamma}$,
$\|\widetilde{Y}_{\gamma}\|_{L^{3}}^{3}\leq C\bigl(1+\|\widetilde{Y}_{\gamma}\|_{L^{4}}^{3}\bigr),$
Then combining with $\|\widetilde{Z}_{\gamma}\|_{L^{\infty}}\lesssim\|\widetilde{Z}_{\gamma}\|_{\mathcal{C}^{\alpha}}$ for every $\alpha\in(1,\frac{1}{2})$, we obtain
$\|\partial_{t}\widetilde{Y}_{\gamma}\|_{H^{-3}}\leq C\Bigl(\|\widetilde{Y}_{\gamma}\|_{H^{1}}+\|\widetilde{Y}_{\gamma}\|_{L^{4}}^{3}+1+\|\widetilde{Z}_{\gamma}\|_{\mathcal{C}^{\alpha}}^{3}+\|R_{\gamma}\|_{H^{-3}}\Bigr).$
Since $\widetilde{Y}_{\gamma}\in L^{2}(0,T;H^{1})\cap L^{4}(0,T;L^{4}),$ we have
$\|\widetilde{Y}_{\gamma}\|_{H^{1}}\in L^{2}(0,T),\qquad\|\widetilde{Y}_{\gamma}\|_{L^{4}}^{3}\in L^{4/3}(0,T).$
Together with $R_{\gamma}\to 0$, this yields the uniform $L^{4/3}(0,T;H^{-3})$-bound.
Consequently, for $0\leq s<t\leq T$,
$\|\widetilde{Y}_{\gamma}(t)-\widetilde{Y}_{\gamma}(s)\|_{H^{-3}}\leq\int_{s}^{t}\|\partial_{t}\widetilde{Y}_{\gamma}(r)\|_{H^{-3}}\,dr\leq|t-s|^{1/4}\,\|\partial_{t}\widetilde{Y}_{\gamma}\|_{L^{4/3}(0,T;H^{-3})}.$
This gives a uniform $1/4$-Hölder-type modulus in time with values in $H^{-3}$, in expectation. ∎
7 Convergence to the stochastic Cahn–Hilliard equation
In this section, we proof the convergence in the main result.
Let $\widetilde{Y}_{\gamma}:=\mathrm{Ext}_{\gamma}Y_{\gamma},\widetilde{Z}_{\gamma}:=\mathrm{Ext}_{\gamma}Z_{\gamma},$ where $Z_{\gamma}$ is the linear stochastic convolution and $Y_{\gamma}$ is the remainder field. We denote $\widetilde{X}_{\gamma}=\widetilde{Y}_{\gamma}+\widetilde{Z}_{\gamma},$ where $\widetilde{Z}_{\gamma}$ is the Fourier extension of the linear stochastic convolution and $\widetilde{Y}_{\gamma}$ is the Fourier extension of the remainder field. Note that the extension operator is uniformly bounded on the discrete Sobolev scales used below, namely
$\|\widetilde{f}_{\gamma}\|_{H^{-1}(\mathbb{T})}\lesssim\|f_{\gamma}\|_{-1,\gamma},\qquad\|\widetilde{f}_{\gamma}\|_{H^{1}(\mathbb{T})}\lesssim\|f_{\gamma}\|_{1,\gamma},\qquad\|\widetilde{f}_{\gamma}\|_{L^{p}(\mathbb{T})}\lesssim\|f_{\gamma}\|_{L^{p}_{\gamma}},$
uniformly in $\gamma$.
Note that $\widetilde{Y}_{\gamma}$ satisfies, in weak form,
$\partial_{t}\widetilde{Y}_{\gamma}=-\nu_{\gamma}\Delta^{2}\widetilde{Y}_{\gamma}-A_{\gamma}\Delta(\widetilde{Y}_{\gamma}+\widetilde{Z}_{\gamma})+\frac{\beta\lambda_{\gamma}}{6}\Delta\bigl((\widetilde{Y}_{\gamma}+\widetilde{Z}_{\gamma})^{3}\bigr)+R_{\gamma},$
where the discrete operators have been transferred to $\mathbb{T}$ through the extension, and the remainders satisfy $R_{\gamma}\rightarrow 0\text{ in }L^{1}(0,T;H^{-3}(\mathbb{T}))$ in probability.
Now we show the convergence of $X_{\gamma}$ to the stochastic Cahn–Hilliard limit.
Theorem 13 (Theorem 1).
Assume that the initial data converge: $\widetilde{X}_{\gamma}(0)\rightarrow X_{0},\text{in }H^{-1}(\mathbb{T});$ Then as $\gamma\to 0$, $\widetilde{X}_{\gamma}$ converges inlaw to $X$ in $L^{2}(0,T;L^{2}(\mathbb{T}))\cap C([0,T],H^{-3}(\mathbb{T})),$ where $X=Y+Z$ is the unique weak solution of the stochastic Cahn–Hilliard equation.
Proof.
We first show the tightness and convergence of $Y_{\gamma}$ via the classical compactness argument.
Since for every $\kappa\in(0,1)$, $H^{1}(\mathbb{T})\hookrightarrow H^{1-\kappa}(\mathbb{T})\hookrightarrow H^{-3}(\mathbb{T}),$ the Aubin–Lions–Simon compactness argument yields relative compactness of $\{\widetilde{Y}_{\gamma}\}_{\gamma}$ in $L^{2}(0,T;L^{2}(\mathbb{T}))$ from the uniform bounds
$\widetilde{Y}_{\gamma}\ \text{bounded in }L^{2}(0,T;H^{1}),\qquad\partial_{t}\widetilde{Y}_{\gamma}\ \text{bounded in }L^{4/3}(0,T;H^{-3}).$
On the other hand, the time Hölder estimate (38) and the uniform $L^{\infty}(0,T;H^{-1})$-bound imply, by the Arzelà–Ascoli criterion in the weak topology together with the compact embedding $H^{-1}(\mathbb{T})\hookrightarrow H^{-3}(\mathbb{T}),$ that the laws of $\{\widetilde{Y}_{\gamma}\}_{\gamma}$ are tight in $C([0,T],H^{-3}(\mathbb{T})).$ Combining the two compactness statements proves the tightness of $\{\widetilde{Y}_{\gamma}\}_{\gamma}$ in
$L^{2}(0,T;H^{1-\kappa}(\mathbb{T}))\cap C([0,T],H^{-3}(\mathbb{T})).$
By Skorokhod’s representation theorem, there exists a probability space $(\Omega,\mathcal{F},\mathbb{P})$ and a limit point
$Y\in L^{\infty}(0,T;H^{-1})\cap L^{2}(0,T;H^{1})\cap L^{4}((0,T)\times\mathbb{T}),$
so that along a subsequence,
$\widetilde{Y}_{\gamma}\to Y\qquad\text{a.s. in }L^{2}(0,T;L^{2}(\mathbb{T}))\cap C([0,T],H^{-3}(\mathbb{T})),$
and $\widetilde{Z}_{\gamma}\to Z$ a.s. in $C([0,T],\mathcal{C}^{\alpha}(\mathbb{T})$. Using the interpolation inequality, the uniform $L^{4}$-bound and strong $L^{2}$-convergence imply that
$\widetilde{X}_{\gamma}=\widetilde{Y}_{\gamma}+\widetilde{Z}_{\gamma}\to Y+Z\qquad\text{strongly in }L^{3}((0,T)\times\mathbb{T}),$
Therefore
[…]
Now test the weak form of the $\widetilde{Y}_{\gamma}$-equation against any $\varphi\in C^{\infty}(\mathbb{T})$, integrate in time from 0 to $t$, and pass to the limit. The linear terms pass to the limit by the strong $L^{2}$-convergence of $\widetilde{Y}_{\gamma}$, the convergence $\mu_{\gamma}\to\mu$, and the convergence of $\widetilde{Z}_{\gamma}$ to $Z$. The cubic term passes to the limit by the $L^{1}$-convergence above. Combining with the convergences of remainder $R_{\gamma}$, and the initial condition, we conclude that the limit $Y$ satisfies the shifted equation
$\partial_{t}Y=-\nu\Delta^{2}Y+\frac{(1+\beta)\lambda}{2}\Delta(Y+Z)+\frac{\beta\lambda}{6}\Delta((Y+Z)^{3}),\quad Y(0)=X_{0},$
in weak sense. By the uniqueness of the shifted equation, then all subsequences have the same limit, and therefore the full sequence converges: $\widetilde{Y}_{\gamma}\Longrightarrow Y$ in $L^{2}(0,T;H^{1-\kappa}(\mathbb{T}))\cap C([0,T],H^{-3}(\mathbb{T})),\forall\kappa\in(0,1).$ Therefore,
$\widetilde{X}_{\gamma}=\widetilde{Y}_{\gamma}+\widetilde{Z}_{\gamma}\Longrightarrow Y+Z=:X$
in $L^{2}(0,T;H^{1-\kappa}(\mathbb{T}))\cap C([0,T],H^{-3}(\mathbb{T})).$ Combining with the weak formulations for $Y$ and $Z$, we obtain that
$dX=\Bigl(-\nu\Delta^{2}X-\frac{(1+\beta)\lambda}{2}\Delta X+\frac{\beta\lambda}{6}\Delta(X^{3})\Bigr)\,dt+\sigma_{*}\partial_{x}dW(t,x),\qquad X(0)=X_{0}.$
Moreover, the uniqueness of stochastic Cahn–Hilliard equation yields that every subsequence of $\{\widetilde{X}_{\gamma}\}_{\gamma}$ has the same limit $X$. This completes the proof. ∎